### v1

# Comparison of Offset and Ratio Weighted Regressions in Tweedie Models with Application to Mid-Term Cancellations

Abstract
In property and casualty insurance, particularly in automobile insurance, risk exposure is traditionally associated with the coverage duration. However, factors such as early contract cancellations demand more precise modelling to ensure accurate premium pricing. This study introduces and compares two approaches for modelling total claims (or loss costs) in insurance portfolios with a high proportion of policies that have partial-year exposure: the offset and ratio methods. We demonstrate that both approaches can be viewed as weighted regressions under the Tweedie distribution framework. Through an analysis based on the financial balance property, we find that the ratio approach outperforms the offset method. This comparison is illustrated using an automobile insurance portfolio, where a significant share of policyholders terminate their contracts before the coverage period concludes.
Keywords Tweedie distribution, weighted regression, offset variable, policy cancellation, stochastic order, positive definite matrix
1 Introduction
In insurance, particularly within the property and casualty sector like automobile insurance, risk exposure has traditionally been associated with the duration of coverage (see Denuit et al. (2007)). Risk exposure refers to the extent to which an insured entity is subject to potential claims or losses over the policy period. For example, under a linear interpretation of exposure, a one-year policy would be considered twice as risky as a six-month policy, assuming identical characteristics for the insured. This concept is central to many actuarial research projects and practical applications.
Given its importance, it becomes essential to develop a comprehensive framework for incorporating various dimensions of risk exposure into the modelling of random variables used for ratemaking. Such models ensure that the premiums charged accurately reflect the underlying risks.
More formally, we are interested in modelling the random variable $Y_{i}$, representing either the total number of claims or the total claims amount (also referred to as the loss cost) for contracts $i=1,\dots,n$, where $n$ denotes the total number of contracts in the portfolio. Assuming that the expected value corresponds to the premium (as discussed in Denuit et al. (2007) and Frees et al. (2014)), the premium can be expressed as:
$\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{T}\beta\right),$ (1.1)
where $\beta$ is a parameter vector to be estimated, and $\mathbf{X_{i}}=(x_{i,0},\dots,x_{i,q})^{T}$ represents the vector of risk characteristics associated with contract $i$ (with $q$ being a positive integer). These risk characteristics—such as age, sex, and the value of the insured goods—are used as covariates in the regression model, allowing the premium to vary based on these factors. The function $g(\cdot)$ serves as a link function that connects the score $\mathbf{X}_{i}^{T}\beta$ to the expected value, similar to the approach in Generalized Linear Models (GLMs) (see Nelder and Wedderburn (1972)).
An analysis of historical motor insurance data reveals significant variability in exposure periods among insured individuals. This variation often occurs when coverage is cancelled before the contract ends due to reasons such as the sale or replacement of a vehicle, theft, or a total loss resulting from an accident. For modelling purposes, insurers typically treat the exposure period as a fraction of the total number of days in a year. This fraction, referred to as risk exposure and denoted by $t$, is a positive value between 0 and 1, reflecting the proportion of the year that the contract was active.
Formally, we consider a portfolio of $n$ insured contracts, indexed by $i$, some of which have a risk exposure, denoted by $t_{i}$, that is less than 1. To account for this variation in exposure when calculating premiums and estimating model parameters, it is necessary to generalize Equation (1.1). This paper examines two approaches within the maximum likelihood estimation framework to address this generalization:
  1. 1.
    The first approach assumes that the exposure $t_{i}$ is proportional to the mean of the conditional distribution $Y_{i}|\mathbf{X_{i}}$, where $f_{Y_{i}}(\cdot)$ represents the density or probability mass function of $Y_{i}|\mathbf{X_{i}}$, and $y_{i}$ denotes the observed loss cost for contract $i$. Under this assumption, the log-likelihood function for estimating the parameter vector $\beta$ is expressed as:
    $\ell(\bm{\beta}|y_{1},\ldots,y_{n})=\sum_{i=1}^{n}\log\left(f_{Y_{i}}(y_{i})\right),$
    where the mean of $Y_{i}$ being linearly proportional to $t_{i}$, such that:
    $\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{T}\beta\right)\times t_{i}.$
    In this framework, $g(\cdot)$ is typically chosen as a logarithmic link function, resulting in:
    $\exp\left(\mathbf{X}_{i}^{T}\beta\right)t_{i}=\exp\left(\mathbf{X}_{i}^{T}\beta+\log(t_{i})\right),$
    where $\log(t_{i})$ acts as an offset variable. This method is thus referred to as the offset approach.
  2. 2.
    The second approach, referred to as ratio approach, focuses on annualizing the random variable $Y_{i}$ to facilitate direct comparisons of policyholders’ claims on an annual basis. In this method, a new random variable $Z_{i}=\frac{Y_{i}}{t_{i}}$ is defined for each contract $i=1,\dots,n$, resulting a normalized set of claim amounts $z_{1},\dots,z_{n}$.
    To distinguish it from the offset approach, the mean of $Z_{i}|\mathbf{X_{i}}$ is assumed to no longer depend on $t_{i}$, and the density or probability mass function of $Z_{i}|\mathbf{X_{i}}$ incorporates the risk exposure through this weighting. The log-likelihood function for estimating $\beta$ is expressed as:
    $\ell(\bm{\beta}|z_{1},\ldots,z_{n})=\sum_{i=1}^{n}\log\left(f_{Z_{i}}(z_{i})\right),$
    where $f_{Z_{i}}(\cdot)$ is the density or probability mass function of $Z_{i}|\mathbf{X_{i}}$, and the mean is given by:
    $\mathrm{E}\!\left[Z_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{T}\beta\right),$
    where $g(\cdot)$ a logarithmic link, the expected loss cost for an insured $i$ exposed to risk over a period of $t_{i}$ years is then expressed:
    $\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=\mathrm{E}\!\left[t_{i}Z_{i}|\mathbf{X_{i}}\right]=\exp\left(\mathbf{X}_{i}^{T}\beta\right)\times t_{i}=\exp\left(\mathbf{X}_{i}^{T}\beta+\log(t_{i})\right),$
    which has the same form as the expected loss cost from the offset approach.
It is important to emphasize that in the offset approach, the random variable $Z_{i}$ can be used to estimate the parameter vector $\beta$, but this method should not be confused with the ratio approach. Similarly, in the ratio approach, the random variable $Y_{i}$ can also be used, but this method should not ne mistaken for the offset approach.
We refer to the offset approach when risk exposure is only incorporated into premium modelling as an offset variable. Conversely, we refer to the ratio approach when risk exposure is only included as a weight in the premium modelling process. These two approaches represent distinct methods for integrating risk exposure, ensuring that premiums are appropriately aligned with the insured’s actual level of risk.
1.1 Examples for claim counts
To clarify the issue, we can illustrate the situation with a preliminary comparison of the two approaches, assuming that $Y_{i}$ represents the number of claims for insured $i$.
Example 1.
Under the ratio approach, we assume that $Z_{i},i=1,\ldots,n$ follows a Poisson distribution with mean $\zeta_{i}=\exp(\mathbf{X}^{T}_{i}\bm{\beta})$. In this framework, it can be shown that the ratio approach corresponds to a Poisson weighted regression and yields the same inference results of the parameter vector $\beta$ as in the offset approach, where $Y_{i},i=1,\ldots,n$ is Poisson distributed with mean $\mu_{i}=t_{i}\exp(\mathbf{X}^{T}_{i}\bm{\beta})=\exp(\mathbf{X}^{T}_{i}\bm{\beta}+\log(t_{i}))$.
Proof.
The log-likelihood function of the offset and ratio approaches are given as follows:
  • Offset Approach:
    $\ell(\bm{\beta}|y_{1},\ldots,y_{n})$ $=\sum_{i=1}^{n}\log\left(\frac{\exp(-\mu_{i})\mu_{i}^{y_{i}}}{y_{i}!}\right)$
    $=\sum_{i=1}^{n}\left(-t_{i}\exp\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)+y_{i}\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)+y_{i}\log(t_{i})-\log(y_{i}!)\right)$
    $\propto\sum_{i=1}^{n}\left(-\exp(\mathbf{X}^{T}_{i}\bm{\beta})+z_{i}\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)\right)\times t_{i}.$
  • Ratio Approach:
    $\ell(\bm{\beta}|z_{1},\ldots,z_{n})$ $=\sum_{i=1}^{n}t_{i}\log\left(\frac{\exp(-\zeta_{i})\zeta_{i}^{z_{i}}}{z_{i}!}\right)$
    $=\sum_{i=1}^{n}\left(-t_{i}\exp\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)+t_{i}z_{i}\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)-\log((t_{i}z_{i})!)\right)$
    $\propto\sum_{i=1}^{n}\left(-\exp(\mathbf{X}^{T}_{i}\bm{\beta})+z_{i}\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)\right)\times t_{i}.$
For the Poisson distribution, the offset and the ratio approaches are equivalent, meaning there is no need to choose between them. ∎
Example 2.
In the offset approach, we assume that $Y_{i},i=1,\ldots,n$ follows a zero-inflated Poisson distribution characterized by the parameters $\phi$ (for the zero-inflation) and $\mu_{i}=\exp(\mathbf{X}^{T}_{i}\bm{\beta}+\log(t_{i}))$ (from the Poisson distribution).
In the ratio approach, we assume that $Z_{i},i=1,\ldots,n$ follows a zero-inflated Poisson distribution characterized by the parameters $\phi$ (for the zero-inflation) and $\zeta_{i}=\exp(\mathbf{X}^{T}_{i}\bm{\beta})$ (from the Poisson distribution).
It can be shown that these two approaches yield different inference results for the estimation of the parameter vector $\beta$.
Proof.
For a portfolio of $n$ contracts, we can develop the log-likelihood of the offset approach to see that both approaches are not equivalent:
$\ell(\bm{\beta}|y_{1},\ldots,y_{n})$ $=\sum_{i=1}^{n}\log\left(I(y_{i}=0)\phi+(1-\phi)\frac{\exp(-\mu_{i})\mu_{i}^{y_{i}}}{y_{i}!}\right)$
$\neq\sum_{i=1}^{n}\log\left(I(z_{i}=0)\phi+(1-\phi)\frac{\exp(-\zeta_{i})\zeta_{i}^{z_{i}}}{z_{i}!}\right)\times t_{i}=\ell(\bm{\beta}|z_{1},\ldots,z_{n}),$
In contrast to the Poisson distribution, a decision must be made between the offset and ratio methods when we want to suppose that the number of claims follows this form of zero-inflated distribution. ∎
Depending on the distribution used for $Y_{i}$, it is not straightforward to determine whether both approaches are similar. Additionally, deciding which approach should be preferred in modelling is not simple. Even if the underlying distribution is the same, the parameters obtained from each approach may differ, which could result in varying premiums for each insured.
1.2 Structure of the paper
This paper emphasizes the differences between the offset and ratio regression approaches in modelling contract loss costs using the Tweedie distribution, rather than focusing on the number of claims. Section 2 introduces the notation and model assumptions, detailing the key properties of the Tweedie distribution and the parameter inference methods employed. Section 3 explains the offset and ratio approaches, as well as the link between them. We show that both approaches can be viewed as weighted regressions, where only the form of the weight differentiates the approaches. An analysis of asymptotic expectations, stochastic dominance, and the gap between aggregate loss costs and estimated premiums using simulated data is presented in Section 4. In Section 5, an empirical illustration of both approaches is provided, based on a sample of automobile insurance data from a Canadian insurer. The last section concludes.
2 GLM with Tweedie-distributed Response Variable
The loss cost of a contract is the total amount of claims, calculated as the aggregate sum of individual claim costs. Formally, let $N$ denote the he total number of claims for a contract, and let $Z_{k}$ represent the cost of the $k$-th claim, where $k,k=1,\ldots,N$ if $N>0$. Thus, assuming $Y$ denotes the loss cost of a contract, as $Y$ is obtained as follows:
$Y=\begin{cases}\sum_{k=1}^{N}Z_{k}&\text{if $N>0$}\\ \\ 0&\text{if $N=0$.}\end{cases}$
Even though individual claim costs $Z_{1},\ldots,Z_{N}$ are continuous random variables, we can demonstrate that the loss cost $Y$ is not a continuous random variable. Specifically, the probability of a null loss cost is non-zero, indicating that the loss cost has a mass at zero and is therefore a semi-continuous random variable. As a result, insurers often use the Tweedie distribution to model a contract premium when the loss cost is the response variable. The rationale for using the Tweedie distribution is its ability to represent a mixture distribution (both continuous and discrete), as noted by Tweedie et al. (1984). Additionally, Delong et al. (2021) provides a comprehensive summary of this mixture distribution, with relevant applications in actuarial science.
Accordingly, we consider the Tweedie distribution as a mixture distribution to model the premium using the loss cost as the response variable. The justification for employing the GLM theory in this context is based on the fact that the Tweedie distribution, as a mixture distribution, belongs to the linear exponential family of distributions, as discussed in Tweedie et al. (1984) and Jørgensen (1987). Before presenting the GLM approach in the context of the Tweedie distribution, we review the Tweedie distribution by referring to the works of Tweedie et al. (1984), Jørgensen (1987), and Delong et al. (2021).
2.1 Exponential Family Form of Tweedie Distribution
Definition 2.1.1.
The density function of the Tweedie distribution as a mixture distribution is given by:
$f(y|\mu,w,\phi)=\exp\left(\frac{w}{\phi}\left(\frac{\mu^{1-p}}{1-p}y-\frac{\mu^{2-p}}{2-p}\right)+a(y,w,\phi)\right),y\in\mathbf{R},$ (2.1)
where:
  • $\mu>0$ denotes the mean parameter;
  • $\phi>0$ denotes the dispersion parameter;
  • $w>0$ denotes the weight parameter;
  • $p\in]1,2[$ denotes the variance parameter;
  • $a(\cdot)$ denotes a function that does not depend on $\mu$.
It is important to note that assuming the variance parameter lies within the interval ]1,2[]1,2[] 1 , 2 [ ensures that the Tweedie distribution density is defined as described above. As noted by Tweedie et al. (1984) and Jørgensen (1987), the Tweedie distribution can be generalized for any positive variance parameter $p>0$. However, they also mention that when the variance parameter lies within the interval ]0,1[]0,1[] 0 , 1 [, the Tweedie distribution does not belong to the exponential family. The density function $f(\cdot)$ mentioned above is expressed in the form of the exponential family density function. In the case of the Tweedie distribution, this family is sometimes referred to as the Exponential Dispersion Family (EDF) due to the inclusion of the dispersion parameter (see Jorgensen (1997)).
It should also be noted that $f(\cdot)$ lacks a specific form because the function $a(\cdot)$ is unknown. To determine $a(.)$, one can refer to Delong et al. (2021), which provides a comprehensive overview of the Tweedie distribution. In this paper, we do not focus on determining $a(\cdot)$ since it does not affect the regression approaches discussed here. Instead, we concentrate on the canonical parameter and the canonical link, which are relevant to insurance pricing. These two quantities for the Tweedie distribution are defined as follows:
  • $\theta$ denotes the canonical parameter and is identified by: $\theta=\frac{\mu^{1-p}}{1-p}$;
  • The canonical link function is defined by: $s>0\mapsto\frac{s^{1-p}}{1-p}$.
It is important to note that the canonical link connects the canonical parameter to the mean parameter of distributions within the exponential family. In car insurance pricing, the canonical link function is often employed within the framework of GLM to ensure the balance property, as described by Nelder and Wedderburn (1972). The balance property ensures that the sum of the estimated premiums equals the sum of the observed loss costs. The rationale behind this equality is that pricing models must replicate the observed risk at each portfolio level as accurately as possible. Consequently, the primary objective for insurers is to develop models that closely reproduce the observed risk across all portfolio levels.
However, when the response variable follows a Tweedie distribution, interpreting the premium through the canonical link becomes challenging. Therefore, as discussed in the following section, choosing the log link to model the premium instead of the canonical link offers a more straightforward interpretation of the premium. It is important to note, however, that using the log link in premium modelling introduces a gap between the sum of the observed loss costs and the sum of the estimated premiums. This paper will analyze this gap in detail.
For the remainder of this paper, we denote the parameters of the Tweedie distribution by $(\mu,w,\phi,p)$ and we define $\mathcal{T}$ as the set of Tweedie distributions $(\mu,w,\phi,p)$ where the variance parameter $p$ lies within the interval ]1,2[]1,2[] 1 , 2 [.
2.2 Notations
We consider $n$ independent insurance contracts. For each contract $i$ (where $i=1,\ldots,n$), the following variables are defined:
  • $Y_{i}$: A random variable representing the loss cost for contract $i$.
  • $y_{i}$: The observed loss cost for contract $i$.
  • $t_{i}$ The risk exposure for contract $i$.
Additionally, we consider $q$ risk factors, denoted by $\bm{x_{1}},\ldots,\bm{x_{q}}$. For each contract $i$, the following are also defined:
  • $x_{i,j}$: The observed characteristic for contract $i$ corresponding to risk factor $\bm{x_{j}}$, where $j=1,\ldots,q$.
  • $\mathbf{X}_{i}$: The vector of risk characteristics associated with contract $i$, defined as $\mathbf{X}_{i}=\left(1,x_{i,1},\ldots,x_{i,q}\right)^{T}$.
Finally, the matrix $\bm{X}$, representing the risk factors, is given by:
$\bm{X}=\begin{pmatrix}1&x_{1,1}&\ldots&x_{1,q}\\ 1&x_{2,1}&\ldots&x_{2,q}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n,1}&\ldots&x_{n,q}\end{pmatrix}$
This matrix $\bm{X}$ captures all the observed risk characteristics across the contracts.
2.3 Assumptions and properties
The main assumptions of GLM are as follows:
  1. 1.
    The distribution of $Y_{i}|\mathbf{X}_{i}$ belongs to the exponential linear family for all contract $i$ where $i=1,\ldots,n$.
  2. 2.
    $Y_{i_{1}}|\mathbf{X}_{i_{1}}\perp\!\!\!\perp Y_{i_{2}}|\mathbf{X}_{i_{2}},\forall i_{1},i_{2}=1,\ldots,n\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ i_{1}\neq i_{2}$.
  3. 3.
    $\bm{X}$ has full rank $q+1$.
The first assumption in the case of the Tweedie distribution is equivalent to stating that $Y_{i}|\mathbf{X}_{i}$ follows a Tweedie distribution with parameters $(\mu_{i},w_{i},\phi,p)$. The second assumption implies the independence of the loss costs between contracts, given the risk characteristics. It is important to note that these first two assumptions are fundamental for constructing the likelihood function. The third assumption is essential for calculating the variance of the estimator of $\mu_{i}$.
The concept of GLM revolves around statistical inference on $\mu_{i}$, which is considered the premium for contract $i$ premium in insurance pricing. This consideration is justified by the fact that the mean parameter $\mu_{i}$, regardless of the distribution of $Y_{i}|\mathbf{X}_{i}$, corresponds to the constant closest to $Y_{i}|\mathbf{X}_{i}$ in terms of Euclidean distance (for further details, see Denuit et al. (2007)).
Thus, for modelling the premium of contract $i$, we assume that the logarithm of this premium ($\mu_{i}$) linearly depends on the characteristics vector $\mathbf{X}_{i}$ as follows:
$\mu_{i}$ $=\exp\left(\bm{X}^{T}_{i}\beta\right)=\exp\left(\beta_{0}\right)\prod_{j=1}^{q}\exp\left(\beta_{j}x_{i,j}\right)=b_{0}\times\prod_{j=1}^{q}r_{j},$ (2.2)
where:
  • $\beta$: The parameter vector, defined as $\beta=(\beta_{0},\beta_{1},\ldots,\beta_{q})^{T}$ where $\beta_{j}\in\mathbf{R}$ for $j=0,1,\ldots,q$;
  • $b_{0}$: The basic premium, defined as $b_{0}=\exp\left(\beta_{0}\right)$.
  • $r_{j}$: The relativity factors associated with each characteristic, defined as $r_{j}=\exp\left(\beta_{j}x_{i,j}\right)$.
The advantage of using the logarithm function as the link function is that the premium $\mu_{i}$ for contract $i$ can be expressed as the product of the basic premium $b_{0}$ and relativity factors $r_{j},j=1,\ldots,q$. While the basic premium $b_{0}$ is consistent across all contracts, the relativity factors depend on the specific risk characteristics of each contract. Consequently, statistical inference on $\mu_{i}$ is equivalent to estimating the parameter vector $\beta$.
To conclude, it is pertinent to introduce an interesting property of the Tweedie distribution, as discussed in Denuit and Trufin (2019), which will be valuable for further developments.
Property 2.3.1 (Tweedie Invariance Property).
If a random variable $Z$ is Tweedie-distributed with parameters $(\mu,w,\phi,p)$ and belongs to the set $\mathcal{T}$, then $tZ$, for any positive $t$, is also Tweedie-distributed with parameters $(t\mu,\frac{w}{t^{2-p}},\phi,p)$ and belongs to the same set $\mathcal{T}$. Conversely, if $tZ$, for any positive $t$, is Tweedie distributed with parameters $(\mu,w,\phi,p)$ and belongs to the set $\mathcal{T}$, then $Z$ is also Tweedie-distributed with parameters$(\frac{\mu}{t},wt^{2-p},\phi,p)$ and belongs to the same set $\mathcal{T}$. Formally, we have:
$Z\in\mathcal{T}\Longleftrightarrow tZ\in\mathcal{T},\forall t>0.$
2.4 Log-likelihood Function
The GLM employs the maximum likelihood approach to estimate the parameter vector $\beta$ (Nelder and Wedderburn, 1972). The maximum likelihood approach maximizes the likelihood function, which depends on the parameter vector $\beta$ given the observations $y_{1},\ldots,y_{n}$. For a detailed discussion on the theory of maximum likelihood, see Casella and Berger (2024). In the case of the Tweedie distribution, by utilizing the conditional independence between contracts and Definition 2.1.1, the likelihood function is expressed as follows:
$L(\beta_{0},\ldots,\beta_{q})=\prod_{i=1}^{n}\exp\left(\frac{w_{i}}{\phi}\left(\frac{\mu^{1-p}}{1-p}_{i}y_{i}-\frac{\mu^{2-p}}{2-p}_{i}\right)+a(y_{i},w_{i},\phi)\right).$
In practice, the maximum likelihood approach often uses the logarithm of the likelihood function as the objective function instead of the likelihood function itself. This is because a function and all strictly increasing transformations of that function reach their maximum at the same point (See Boyd and Vandenberghe (2004) for the theory on function optimization). Therefore, in the Tweedie case, the logarithm of the likelihood function (commonly referred to as the log-likelihood function) is given by (2.3).
$\ell(\beta_{0},\ldots,\beta_{q})$ $=\sum_{i=1}^{n}\left(\frac{w_{i}}{\phi}\left(\frac{\mu^{1-p}_{i}}{1-p}y_{i}-\frac{\mu^{2-p}_{i}}{2-p}\right)+a\left(y_{i},w_{i},\phi,p\right)\right)$ (2.3)
$\propto\frac{1}{\phi}\sum_{i=1}^{n}w_{i}\left(\frac{\mu^{1-p}_{i}}{1-p}y_{i}-\frac{\mu^{2-p}_{i}}{2-p}\right).$ (2.4)
Additionally, considering equation (2.4), we note that maximizing the log-likelihood in the Tweedie case is equivalent to performing a weighted regression, where each contract $i$ is weighted by $w_{i}$. The log-likelihood maximization can also be interpreted as finding the zeros of the gradient function associated with the model. The gradient function is the vector of the first partial derivatives of the log-likelihood function ($\frac{\partial\ell(\beta_{0},\ldots,\beta_{q})}{\partial\beta_{j}},j=0,\ldots,q$). For the Tweedie case, the gradient function is derived as follows:
$\nabla(\beta)$ $=\left(\frac{\partial\ell(\beta_{0},\ldots,\beta_{q})}{\partial\mu_{i}}\frac{\partial\mu_{i}}{\partial\beta_{0}},\ldots,\frac{\partial\ell(\beta_{0},\ldots,\beta_{q})}{\partial\mu_{i}}\frac{\partial\mu_{i}}{\partial\beta_{j}}\right)^{t}$
$=\frac{1}{\phi}\left(\sum_{i=1}^{n}w_{i}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}},\ldots,\sum_{i=1}^{n}w_{i}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}}x_{i,q}\right)^{t}$
$=\frac{1}{\phi}\bm{X}^{t}\bm{D}\bm{R};$
The vector $\bm{R}$ and the matrix $\bm{D}$ are defined as follows:
$\bm{R}=\left(\frac{y_{i}}{\mu_{i}}-1\right)_{i=1,...,n}\ \bm{D}=\operatorname{diag}\left(w_{i}\mu^{2-p}_{i}\right)_{i=1,...,n}.$ (2.5)
2.4.1 Iterative Algorithm
Generally, GLM theory adopts the IRLS (Iteratively Reweighted Least Squares) algorithm for maximizing the log-likelihood, which is a variant of the Newton-Raphson algorithm (For a detailed discussion on the Newton-Raphson algorithm, refer to Boyd and Vandenberghe (2004)). It is important to note that the IRLS algorithm is based on the Fisher Information Matrix, the theory of which can be found in Casella and Berger (2024). In this section, we review the Fisher Information Matrix in the context of the Tweedie distribution and justify its invertibility. We also recall the definition of positive definite matrices and their properties, as discussed by Horn and Johnson (2012).
Let $\beta_{(k)}$ denote the estimate of $\beta$ at iteration $k$, The IRLS algorithm is defined as follows:
$\beta_{(k+1)}=\beta_{(k)}+\left(I_{n}\left(\beta_{(k)}\right)\right)^{-1}\nabla\left({\beta_{(k)}}\right);$ (2.6)
$I_{n}(\cdot)$ is the Fisher information matrix and is defined as follows:
$I_{n}(\beta)$ $=-\mathrm{E}\!\left[\text{Hess}\left({\beta}\right)\right]$
$=\frac{1}{\phi}\left(\sum_{i=1}^{n}w_{i}\mu^{2-p}_{i}x_{i,j_{1}}x_{i,j_{2}}\right)_{{}_{j_{1},j_{2}=0,\ldots,q}}$
$=\frac{1}{\phi}\bm{X}^{t}\bm{D}\bm{X},$
where $\text{Hess}\left({\beta}\right)$ is the Hessian matrix. Referring to equation (2.6), we note that the IRLS algorithm relies on the invertibility of the Fisher Information Matrix. The inverse of the Fisher Information Matrix is also used to calculate the asymptotic covariance matrix of the parameter vector estimator. As demonstrated in this paper, comparing the covariance matrices of the parameter vector estimators is crucial for comparing premium estimators in the offset and ratio approaches. Therefore, analyzing the invertibility of the Fisher Information Matrix is of prime importance.
We now review the properties of positive definite matrices and introduce the following notations:
  • $\mathcal{M}_{n,m}$ denotes the set of matrices with $n$ rows and $m$ columns.
  • $\mathcal{P}_{m}$ denotes the set of positive definite matrices.
Definition 2.4.1.
A matrix $M\in\mathcal{M}_{m,m}$ is positive definite ($M\succ 0$), if the following condition holds:
$x^{t}Mx>0;\forall x\in\mathbb{R}^{m}.$
Property 2.4.1.
With $\mathcal{M}_{n,m}$ and $\mathcal{P}_{m}$ defined above, the following relationships can be stated:
$A\in\mathcal{P}_{m}$ $\Longleftrightarrow A^{-1}\in\mathcal{P}_{m};$ (2.7)
$\forall A,B\in\mathcal{P}_{m}:A-B\succ 0$ $\Longleftrightarrow B^{-1}-A^{-1}\succ 0;$ (2.8)
$\forall A\in\mathcal{P}_{m},\forall X\in\mathcal{M}_{m,q}:X^{t}AX\succ 0$ $\Longleftrightarrow X\leavevmode\nobreak\ \text{has rank}\leavevmode\nobreak\ q.$ (2.9)
Property (2.7) serves as a necessary and sufficient condition for proving the invertibility of a matrix. Thus, to prove the existence of the inverse of the Fisher Information Matrix, we must show that this matrix is positive definite. Using the GLM theory assumption regarding the rank of the matrix $\bm{X}$ and the necessary and sufficient condition given in (2.9), we conclude that the Fisher Information Matrix is indeed positive definite. The necessary and sufficient condition given in (2.8) is also helpful when comparing premium estimators in the offset and ratio approaches.
2.4.2 Remarks
The variance and weight parameters exclusively influence the estimates of the parameter vector in the IRLS algorithm. However, as stated in equation (2.10), the dispersion parameter does not affect these estimates, and therefore, it is not a focus of this paper. Finally, if the weight of each contract is known and the variance parameter is also known, the GLM in the Tweedie case involves using the IRLS algorithm to obtain an estimate of the parameters vector.
$\left(I_{n}(\beta)\right)^{-1}\nabla\left(\beta\right)=\left(X^{t}DX\right)^{-1}X^{t}DR.$ (2.10)
It is also noteworthy that the IRLS algorithm is implemented in the R software through the glm function.
3 The Offset and the Ratio Approaches
To present the offset and ratio regression approaches within the context of a Tweedie distribution and to examine their similarities and differences, we continue to consider $n$ independent contracts, along with the same notations and assumptions as in the previous section. The primary distinction from the previous section is that, in these two regression approaches, the premium of contract $i$ is adjusted to account for its risk exposure $t_{i}$.
3.1 Offset Approach
As mentioned in Section 1.1, the statistical inference on the parameter vector $\beta$ in offset approach is performed with the lost costs $Y_{1},\ldots,Y_{n}$. To account for the logarithm of a contract’s risk exposure as the offset variable, we assume that $Y_{i}|\mathbf{X}_{i}$ follows the Tweedie distribution with parameters $(\mu_{i},1,\phi,p)$ where $\mu_{i}$ is defined as $\mu_{i}=t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right)$. We also denote an estimate of $\mu_{i}$ in the offset regression by $\widehat{\mu}^{O}_{i}$, obtained as follows:
$\widehat{\mu}^{O}_{i}=\exp\left(\mathbf{X}^{T}_{i}\widehat{\beta}^{O}+\log(t_{i})\right),$ (3.1)
where $\widehat{\beta}^{O}$, an estimate of $\beta$ in offset approach, is obtained using the IRLS algorithm with the following quantities:
  • Log-likelihood Function:
    $\ell^{O}(\beta_{0},\ldots,\beta_{q})=\sum_{i=1}^{n}\left(\frac{1}{\phi}\left(\frac{\mu^{1-p}_{i}}{1-p}y_{i}-\frac{\mu^{2-p}_{i}}{2-p}\right)+a^{O}\left(y_{i},\phi,p\right)\right).$ (3.2)
  • Gradient Function:
    $\nabla^{O}(\beta)$ $=\frac{1}{\phi}\left(\sum_{i=1}^{n}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}},\ldots,\sum_{i=1}^{n}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}}x_{i,q}\right)^{t}$ (3.3)
    $=\frac{1}{\phi}\bm{X}^{t}\bm{D}^{O}\bm{R}^{O};$ (3.4)
    The vector $\bm{R}^{O}$ and the matrix $\bm{D}^{O}$ are defined as follows:
    $\bm{D}^{O}=\operatorname{diag}\left(t^{2-p}_{i}\exp\left((2-p)\mathbf{X}^{T}_{i}\beta\right)\right)_{i=1,...,n},\bm{R}^{O}=\left(\frac{y_{i}}{\mu_{i}}-1\right)_{i=1,...,n}.$
3.2 Ratio Approach
We saw in Section 1.1 that the ratio-based method can be interpreted as a form of weighted regression. In this framework, the response variable is defined as the ratio of the contract’s loss cost to its risk exposure, with each contract being assigned a weight proportional to its respective risk exposure. We denote the ratio of contract $i$’s loss cost $Y_{i}$ and its risk exposure $t_{i}$ by $Z_{i}$, defined as $Z_{i}=\frac{Y_{i}}{t_{i}}$, for $i=1,\ldots,n$. To account for the risk exposure $t_{i}$ as the weight of contract $i$, we also assume $Z_{i}|\mathbf{X}_{i}$ follows a Tweedie distribution with parameters $(\zeta_{i},t_{i},\phi,p)$, where the mean parameter $\zeta_{i}$ is assumed to be:
$\zeta_{i}=\exp\left(\mathbf{X}^{T}_{i}\beta\right),$ (3.5)
and $\widehat{\zeta_{i}}$ is an estimate of $\zeta_{i}$, equal to $\widehat{\zeta_{i}}=\exp\left(\mathbf{X}^{T}_{i}\widehat{\beta}^{R}\right)$. The vector $\widehat{\beta}^{R}$, an estimate of $\beta$ in ratio approach, is obtained using the IRLS algorithm with the following quantities:
  • Log-likelihood Function: This function is constructed with the observations $z_{1},\ldots,z_{n}$ where $\zeta_{i}=\frac{y_{i}}{t_{i}},i=1,\ldots,n$, as follows:
    $\ell^{R}(\beta_{0},\ldots,\beta_{q})=\sum_{i=1}^{n}\left\{\frac{t_{i}}{\phi}\left(\frac{\zeta^{1-p}_{i}}{1-p}z_{i}-\frac{\zeta^{2-p}_{i}}{2-p}\right)+a^{R}\left(z_{i},\phi,p\right)\right\}.$ (3.6)
  • Gradient Function:
    $\nabla^{R}(\beta)$ $=\frac{1}{\phi}\left(\sum_{i=1}^{n}t_{i}\frac{z_{i}-\zeta_{i}}{\zeta^{p-1}_{i}},\ldots,\sum_{i=1}^{n}t_{i}\frac{z_{i}-\zeta_{i}}{\zeta^{p-1}_{i}}x_{i,q}\right)^{t}$ (3.7)
    $=\frac{1}{\phi}\bm{X}^{t}\bm{D}^{R}\bm{R}^{R};$ (3.8)
    The vector $\bm{R}^{R}$ and the matrix $\bm{D}^{R}$ are defined as follows:
    $\bm{D}^{R}=\operatorname{diag}\left(t_{i}\exp\left\{(2-p)\mathbf{X}^{T}_{i}\beta\right\}\right)_{i=1,...,n},\bm{R}^{R}=\left(\frac{z_{i}}{\zeta_{i}}-1\right)=\left(\frac{y_{i}}{\mu_{i}}-1\right)_{i=1,...,n},$
    where $\mu_{i}=t_{i}\zeta_{i}=t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right).$
3.3 Link between Offset and Ratio Approaches
Building on the Tweedie Invariance Property defined in Proposition 2.3.1, the equivalence of the two approaches can be demonstrated as follows:
  1. 1.
    The offset approach viewed in the form of the ratio approach
    In the offset approach, $Y_{i}|\mathbf{X}^{T}_{i}$ follows a Tweedie distribution with parameters $(t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)$. We can examine the ratio of the loss cost to the risk exposure $t_{i}$, denoted as $Z_{i}=\frac{Y_{i}}{t_{i}}$, and apply the Tweedie Invariance Property. It can then be demonstrated that $Z_{i}|\mathbf{X}^{T}_{i}$ follows a Tweedie distribution with parameters $(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{2-p},\phi,p)$. More formally, we have the following equivalence:
    $Y|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)\equiv Z|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{2-p},\phi,p).$
    Using equation (2.5) for the calculation of the matrix $\bm{D}$, we can demonstrate that the two Tweedie distributions introduced earlier share the same expression:
    $\bm{D}=\operatorname{diag}\left(t^{2-p}_{i}\exp\left((2-p)\mathbf{X}^{T}_{i}\beta\right)\right)_{i=1,\ldots,n}\equiv\bm{D}^{O}$
    This result provides an additional verification that these two regression approaches are equivalent, leading to identical inference outcomes for the parameter vector $\beta$.
  2. 2.
    The ratio approach viewed in the form of the offset approach
    Similarly, in the ratio approach, where $Z_{i}|\mathbf{X}^{T}_{i}$ is assumed to follow a Tweedie distribution with parameters $(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)$, we can apply the Tweedie Invariance Property to demonstrate that $Y_{i}|\mathbf{X}^{T}_{i}$ follows a Tweedie distribution with parameters $(t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{p-1},\phi,p)$. This equivalence is expressed as:
    $Z|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)\equiv Y|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{p-1},\phi,p).$
    For each form of the Tweedie distributions, we also have the following equivalence for the matrix $\bm{D}$:
    $\bm{D}=\operatorname{diag}\left(t_{i}\exp\left((2-p)\mathbf{X}^{T}_{i}\beta\right)\right)_{i=1,\ldots,n}\equiv\bm{D}^{R}.$
The previous results facilitate a direct comparison of the two approaches. By focusing on the random variable $Z$, the choice is between the following two distributions:
$\text{Tweedie}(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{2-p},\phi,p)\ \ \ \text{{versus}}\ \ \ \text{Tweedie}(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)$
We observe that the choice of approach depends solely on the form of the weights $w_{i}$, as all other parameters remain unchanged. In other words, selecting between an offset or a ratio approach implies working with normalized loss costs, i.e., the random variable $Z_{i}=\frac{Y_{i}}{t_{i}}$, and deciding, for observation $i$, whether to apply a weight of $w_{i}^{O}=t_{i}^{2-p}$ or a weight of $w_{i}^{R}=t_{i}$.
3.4 Weights Analysis
Since the distinction between the two approaches primarily hinges on the choice of the weight parameter $w_{i}$, it is crucial to analyze its impact on modelling loss costs with the Tweedie distribution in greater detail. An initial evident result is the dominance of the weight in the ratio approach compared to that in the offset approach, given by:
$w_{i}^{O}=t_{i}^{2-p}\geq t_{i}=w_{i}^{R},\quad\forall p\in]1,2[,\,t\in]0,1].$
In Figure 1, we analyze in more detail the differences between the weights used in the offset and ratio approaches. The diagonal dashed line represents the weights in the ratio approach $(w_{i}^{R}=t_{i})$, while the other coloured lines illustrate the weights $t_{i}^{2-p}$ employed in the offset approach for various values of the variance parameter $p$.
Figure 1: Comparison of the Weight Parameters in the Ratio and Offset Approaches for Different Risk Exposures ($t$)
For a given value of $p$, we observe that the difference in weights between the offset and ratio approaches diminishes as the risk exposure increases. This indicates that when the contract exposure period is short, the disparity between the weights in both approaches is significant compared to a contract with an exposure period close to one year. In the case where the exposure period equals one year $(t=1)$, the weights in each approach equal 1, which means that these two regression approaches are equivalent. This equivalence between the methods is also evident in the equality between $\bm{D}^{R}$ and $\bm{D}^{O}$ when $t=1$:
$\bm{D}^{R}=\bm{D}^{O}=\operatorname{diag}\left(\exp\left((2-p)\mathbf{X}^{T}_{i}\beta\right)\right)_{i=1,...,n}.$
We can also interpret Figure 1 by varying the variance parameter $p$. As $p$ decreases and approaches 1, the difference between the weights in the offset and ratio approaches becomes negligible. Since the Tweedie distribution converges to the Poisson distribution as $p\rightarrow 1$, it is noteworthy that this observation aligns with the equivalence of the methods, corresponding to the result obtained regarding claim numbers in Example 1 presented at the beginning of the paper. Conversely, as $p$ approaches 2, the disparity in weights between the two approaches becomes more pronounced, indicating that the choice of approach has a more significant impact. Notably, when $p\rightarrow 2$, the Tweedie distribution converges to a gamma distribution, which has historically been employed to model the severity of claims in actuarial science (see Denuit et al. (2007)). In the context of severity modelling, it is well known that the risk exposure $t_{i}$ should not be utilized, implying that the weight should remain constant regardless of the value of the risk exposure $t_{i}$. The ratio approach with $p=1.999$ tends to demonstrate this characteristic. Thus, overall, these results suggest a preference for the ratio approach, which appears appropriate for any value of $p\in]1,2[$.
4 Comparison of the Approaches
According to our analysis presented in Section 3.3, for a given value of the variance parameter $p$, the difference between weights $w_{i}^{O}$ and $w_{i}^{R}$, of the offset and ratio approaches respectively, depends on the risk exposure $t_{i}$. This difference is significant for contracts with shorter exposure periods, while it becomes negligible for those with exposure periods close to one year. However, a limitation of this analysis is that it does not conclusively determine which approach may be more appropriate for insurance pricing.
Our strategy to compare approaches involves utilizing the distribution of the maximum likelihood estimators of the parameter vectors in the offset and ratio approaches to compare the premium estimators [^1] We will use the expression premium estimator instead of estimated premium throughout this discussion, as the estimated premium refers to the value of a single observation derived from the premium estimator, which is a more general mathematical expression. in these two regression methods. We begin by comparing the asymptotic expected values and variances of the premium estimators in both regression approaches. This comparison utilizes the positive definite matrix property described in equation (2.8), which enables a direct evaluation of the asymptotic covariance matrices of the parameter vector estimators in the offset and ratio regression approaches. We conclude our analysis by applying the concept of stochastic order (or stochastic dominance), as discussed in Shaked and Shanthikumar (2007).
4.1 Notations and assumptions
We continue to consider $n$ independent contracts, using the same notations and assumptions as in previous sections. To facilitate comparison, we define the response variable consistently across both approaches:
  • Offset approach: $Z|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(\zeta_{i},t^{2-p}_{i},\phi,p)$ for $i=1,\ldots,n$,
  • Ratio approach: $Z|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(\zeta_{i},t_{i},\phi,p)$ for $i=1,\ldots,n$,
where $\zeta_{i}$ is the premium for the $i$-th contract without adjustment for risk exposure and is linked to the score $\mathbf{X}^{T}_{i}\beta$ as specified in Equation (4.1):
$\zeta_{i}=\exp\left(\mathbf{X}^{T}_{i}\beta\right).$ (4.1)
Using this formulation, the premium estimators for $\zeta_{i}$ in the offset and ratio approaches, denoted by $\bm{\widehat{\zeta}^{O}}_{i}$ and $\bm{\widehat{\zeta}^{R}}_{i}$, are obtained as follows:
$\bm{\widehat{\zeta}^{O}}_{i}$ $=\exp\left(\mathbf{X}^{T}_{i}\bm{\widehat{\beta}^{O}}\right);$ (4.2)
$\bm{\widehat{\zeta}^{R}}_{i}$ $=\exp\left(\mathbf{X}^{T}_{i}\bm{\widehat{\beta}^{R}}\right);$ (4.3)
Where $\bm{\widehat{\beta}^{O}}$ and $\bm{\widehat{\beta}^{R}}$ are the estimators of $\beta$ in the offset and ratio approaches, respectively.
4.2 Expected Values and Variances Comparison of Premium Estimators
To compare the offset and ratio regression approaches, a valuable direction is to examine the entire distribution of the premium estimator for each approach. However, comparing the expected values and variances of the premium estimators may be sufficient for this analysis. Indeed, when evaluating competing estimators, the optimal choice is generally the one with the smallest bias and variance. Therefore, our quantities of interest are the expected values and variances of the premium estimators for both approaches, as outlined in the following propositions.
Proposition 1.
The expected value and variance of premium estimator in offset and ratio approach are obtained as follows:
$\mathrm{E}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]$ $=\exp\left\{\mathbf{X}^{T}_{i}\beta+\frac{\phi}{2}\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}\mathbf{X}_{i}\right\};$ (4.4)
$\mathrm{E}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right]$ $=\exp\left\{\mathbf{X}^{T}_{i}\beta+\frac{\phi}{2}\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\mathbf{X}_{i}\right\};$ (4.5)
$\mathrm{Var}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]$ $=\left(\exp\left\{\phi\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}\mathbf{X}_{i}\right\}-1\right)\left(\mathrm{E}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]\right)^{2};$ (4.6)
$\mathrm{Var}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right]$ $=\left(\exp\left\{\phi\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\mathbf{X}_{i}\right\}-1\right)\left(\mathrm{E}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right]\right)^{2}.$ (4.7)
Proof.
The key to the proof is to note that, for a random variable $Z$, the relationship $\log(Z)\sim\mathcal{N}(\mu,\sigma^{2})$ holds. The expected value and variance of $Z$ are then given by:
  • $\mathrm{E}\!\left[Z\right]=\exp\{\mu+\frac{1}{2}\sigma^{2}\}$;
  • $\mathrm{Var}\!\left[Z\right]=\left(\exp\left\{\sigma^{2}\right\}-1\right)\left(\mathrm{E}\!\left[Z\right]\right)^{2}$.
Additionally, the asymptotic properties of the maximum likelihood estimator (MLE) allow us to derive the distributions of the parameter vector estimators as follows:
$\bm{\widehat{\beta}^{O}}\sim\mathcal{N}\left(\beta,\phi\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}\right)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \bm{\widehat{\beta}^{R}}\sim\mathcal{N}\left(\beta,\phi\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\right).$
So, using these asymptotic distributions and Equations (4.2) and (4.2), we obtain the following relationships:
$\log\left(\bm{\widehat{\zeta}^{O}}_{i}\right)\sim\mathcal{N}\left(\mathbf{X}^{T}_{i}\beta,\phi\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}\mathbf{X}_{i}\right)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \log\left(\bm{\widehat{\zeta}^{R}}_{i}\right)\sim\mathcal{N}\left(\mathbf{X}^{T}_{i}\beta,\phi\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\mathbf{X}_{i}\right).$
This demonstrates how the expected values and variances of the premium estimators can be derived using the distributions of the parameter vector estimators. ∎
Proposition 2.
The expected value of the premium estimator in offset approach is smaller than in the ratio approach. Additionally, the variance of the premium estimator in the offset is also smaller than in the ratio approach. Formally, for contract $i$, we obtain:
$\mathrm{E}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]$ $<\mathrm{E}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right];$ (4.8)
$\mathrm{Var}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]$ $<\mathrm{Var}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right].$ (4.9)
Proof.
Referring to the expressions for the expected values and variances of premium estimators given in Proposition 1, we observe that the key difference between these metrics arises from the asymptotic covariance matrices of the parameter vector estimators in the offset and ratio regressions: $\phi\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1},\phi\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}$.
This proof, therefore, focuses on comparing these covariance matrices between the two approaches. As noted in Section 2.4.1, both covariance matrices are well defined and positive definite:
$\phi\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1},\phi\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\in\mathcal{P}_{q+1}.$
Let us define the matrix $M$ as the difference between these two covariance matrices:
$M=\phi\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}-\phi\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}.$
The goal is to show that $M$ is a positive definite matrix. To do so, we can utilize the positive definite matrix property described in Section 2.8. Therefore, proving that $D^{O}-D^{R}$ is a positive definite matrix is sufficient to show that $M$ is also one.
By examining Figure 1, we observe that $D^{O}-D^{R}$ is indeed a positive definite matrix. This can be formally expressed as:
$\forall v=(v_{1},\ldots,v_{n})^{t}\in\mathbf{R}^{n}:v^{t}\left(D^{O}-D^{R}\right)v=\sum_{i=1}^{n}\left(t^{2-p}_{i}-t_{i}\right)\zeta^{2-p}_{i}v^{2}_{i}>0.$
Additionally, using equation (2.9), we note that the following matrix is positive definite because the matrix $\bm{X}$ has full rank $q+1$:
$\bm{X}^{t}\left(\bm{D}^{O}-\bm{D}^{R}\right)\bm{X}=\bm{X}^{t}\bm{D}^{O}\bm{X}-\bm{X}^{t}\bm{D}^{R}\bm{X}.$
Thus $M$ being positive definite implies that the expected value and variance of the premium estimator in the offset approach are smaller than in the ratio approach, as stated in Proposition 2. ∎
From equations (4.4) and (4.5), it is evident that there is a positive bias in the premium estimator for both regression approaches, as $\mathrm{E}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]-\zeta_{i}>0$ and $\mathrm{E}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right]-\zeta_{i}>0$. However, using Proposition 2, we observe that the bias of the premium estimator in the ratio approach is consistently larger than that in the offset approach. Therefore, based on the bias and variance criterion, the premium estimator from the offset approach is superior to that of the ratio approach.
It is important to note that, in insurance pricing, the primary focus is on the bias of the premium estimator. Indeed, the goal for an insurer is to develop a pricing model that significantly minimizes premium estimator bias. Auto-calibration is generally used for this purpose. For further details on premium estimator bias reduction, recent works by Denuit et al. (2021), Lindholm et al. (2023), Denuit et al. (2024) and Denuit and Trufin (2024) can be consulted. However, this paper prioritizes bias reduction at the portfolio level rather than at the individual premium level.
4.3 Random Gaps Comparison
Another comparison between the two approaches can be made based on what we will refer to as the random gaps. More formally, this refers to the following values for the offset and ratio approaches, respectively:
$\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right).$
These expressions represent the difference between two random variables: one corresponding to the loss cost of insured individual $i$, and the other to the premium estimator for the same individual. We refer to the expectation of the random gap as the "portfolio bias". According to Denuit et al. (2024), financial equilibrium at the portfolio level corresponds to a zero portfolio bias. In other words, the expectation of the random gap should be close to zero to ensure financial equilibrium at the portfolio level.
However, due to the positive bias in the premium estimators for both regression approaches, financial equilibrium is not achieved in either the offset or ratio approach at the portfolio level. Using Proposition 2, we note that the portfolio bias in the ratio approach is smaller than in the offset approach. Formally, we establish the following relationship:
$\mathrm{E}\!\left[\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\right]\leq\mathrm{E}\!\left[\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\right].$ (4.10)
This relationship implies that the premium estimator from the ratio approach is more effective than that from the offset approach in achieving financial equilibrium across the entire portfolio. Our objective is to provide a probabilistic framework for comparing the full distributions of these random gaps using stochastic dominance. Specifically, first-order stochastic dominance allows us to demonstrate that the ratio approach results in smaller gaps with higher probability than the offset approach.
Before presenting our results, we first formally define stochastic dominance between two random variables, based on the work of Shaked and Shanthikumar (2007).
Definition 4.3.1.
Consider two random variables, $R_{1}$ and $R_{2}$. We say $R_{2}$ dominates $R_{1}$, denoted by $R_{1}\leq_{st}R_{2}$, if the following condition is satisfied:
$\mathrm{Pr}\!\left(R_{1}>x\right)\leq\mathrm{Pr}\!\left(R_{2}>x\right),\forall x\in\mathbf{R}.$
Additionally, we provide a necessary and sufficient condition for ordering two random variables, as outlined in Shaked and Shanthikumar (2007).
Property 4.3.1.
For two random variables, R1 and R2, the following relationship holds:
$R_{1}\leq_{st}R_{2}\Longleftrightarrow\mathrm{E}\!\left[f(R_{1})\right]\leq\mathrm{E}\!\left[f(R_{2})\right],$
where $f(\cdot)$ is an increasing function for which the expectations exist.
The two previous results enable us to state the following property:
Proposition 3.
The difference between the aggregate loss costs and the sum of the premium estimators is more likely to take a significant value in the offset regression approach than in the ratio regression approach. Formally, for contract $i$, we establish the following stochastic dominance:
$\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\leq_{st}\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right).$ (4.11)
Proof.
The key to this proof is using the Taylor series approximation as discussed in Van der Vaart (2000). Specifically, the idea is to approximate the expected value of a function of a random variable by applying the function to the expected value of the random variable itself.
Consider a function $f(\cdot)$. By applying the first-order Taylor approximation to the function $f(\cdot)$ at the sums of the difference between the aggregate loss cost and the sum of the premium estimators, $\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)$ and $\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)$, we obtain the following results for each contract $i$:
$\mathrm{E}\!\left[f\left(\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\right)\right]$ $\approx f\left(\mathrm{E}\!\left[\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\right]\right);$ (4.12)
$\mathrm{E}\!\left[f\left(\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\right)\right]$ $\approx f\left(\mathrm{E}\!\left[\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\right]\right).$ (4.13)
Using Equation (4.10) and considering $f(.)$ as an increasing function, we conclude that:
$\mathrm{E}\!\left[f\left(t_{i}\sum_{i=1}^{n}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\right)\right]\leq\mathrm{E}\!\left[f\left(t_{i}\sum_{i=1}^{n}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\right)\right].$
This proof shows that the gap between the aggregate loss costs and the sum of the premium estimators is more likely to be larger in the offset regression approach than in the ratio regression approach.
4.3.1 Remarks
In general, for each contract $i,i=1,\ldots,n$, when the premium, $\zeta_{i}$, is linked to the score $\mathbf{X}^{T}_{i}\beta$ as $\zeta_{i}=g^{-1}\left(\mathbf{X}^{T}_{i}\beta\right)$, $\widehat{\beta}$ serves as the estimator of $\beta$ in both the offset and ratio approaches. However, except in the case of the identity link function, $s>0\mapsto g(s)=s$, the premium estimator $\zeta_{i}$ has non-zero asymptotic bias:
$\mathrm{E}\!\left[g^{-1}\left(\mathbf{X}^{T}_{i}\widehat{\beta}\right)\right]\neq g^{-1}\left(\mathbf{X}^{T}_{i}\beta\right).$
Consequently, financial equilibrium based on the expected random gap cannot be achieved unless the identity link is used in premium modelling. For this reason, in GLM, insurers consider the balance property to ensure an alternative financial equilibrium based on the alignment of observed loss costs with estimated premiums (see Nelder and Wedderburn (1972) for a discussion on the relationship between the sum of the predicted values and the sum of the observed values in GLM theory). As discussed in Section 2, this financial equilibrium ensures that the sum of the estimated premiums equals the aggregate observed loss costs across the entire portfolio, as well as at each level of risk factors used in premium modelling.
It is important to note that even if the balance property is not verified in either approach, Proposition 3 indicates that it is more likely for the random gaps to take on larger values in the offset approach than in the ratio approach. However, this alone is insufficient to determine which approach—offset or ratio—is preferable for achieving financial equilibrium under the balance property. Therefore, we need to examine the difference between aggregate observed loss costs and the sum of estimated premiums for both approaches.
4.4 Observed Gaps Comparison
We refer to the observed gaps in each regression approach as defined below:
$\sum_{i=1}^{n}t_{i}\left(z_{i}-\widehat{\zeta}^{R}_{i}\right)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \sum_{i=1}^{n}t_{i}\left(z_{i}-\widehat{\zeta}^{O}_{i}\right),$ (4.14)
where $\widehat{\zeta}^{R}_{i}$ and $\widehat{\zeta}^{O}_{i}$ represent the estimated of $\zeta$ from the ratio and offset approaches, respectively. These values, $\widehat{\zeta}^{R}_{i}$ and $\widehat{\zeta}^{O}_{i}$ can also be viewed as realizations of the premium estimators in the ratio and offset approaches.
Our objective is to demonstrate that the observed gaps are equal to zero in the ratio approach when applied to a homogeneous portfolio. However, under the same conditions, alignment between aggregate loss costs and the sum of estimated premiums does not occur in the offset approach. For a heterogeneous portfolio, the observed gaps in both the offset and ratio approaches are not exactly zero due to the influence of the log-link function in premium modeling. Nonetheless, the relationship between the sum of estimated premiums and the aggregate loss costs in the ratio approach can be expressed mathematically. This formulation offers valuable insights into the characteristics and magnitude of the observed gap in this approach.
Our strategy for examining these gaps involves analyzing the MLE for each approach under both homogeneous and heterogeneous portfolio across contracts. By maintaining the assumption of $n$ independent contracts and applying consistent notation and assumptions from previous sections, we ensure coherence across models and approaches.
4.4.1 Homogeneous Portfolio
First, we will assume the analysis of a homogeneous portfolio for which no explanatory variable is used in the mean parameter of the two approaches under study. More formally, we assume that the loss costs $Z_{1},\ldots,Z_{n}$ are independent and have a common mean parameter $\zeta$. We also assume that $\zeta_{i}=\zeta$ for all contracts $i=1,\ldots,n$. This allows us to express the log-likelihood function for $\zeta$ as follows:
$\ell(\zeta)=\frac{1}{\phi}\left(\frac{\zeta^{1-p}}{1-p}\sum_{i=1}^{n}w_{i}z_{i}-\frac{\zeta^{2-p}}{2-p}\sum_{i=1}^{n}w_{i}\right)+\sum_{i=1}^{n}a\left(z_{i},w_{i},\phi,p\right).$ (4.15)
The first derivative of this log-likelihood function is given by:
$\ell^{{}^{\prime}}(\zeta)=\frac{1}{\phi\zeta^{p}}\left(\sum_{i=1}^{n}w_{i}z_{i}-\zeta\sum_{i=1}^{n}w_{i}\right).$ (4.16)
Thus, the MLE of $\zeta$ , denoted by $\widehat{\zeta}$, is obtained as follows: $\widehat{\zeta}=\frac{\sum_{i=1}^{n}w_{i}z_{i}}{\sum_{i=1}^{n}w_{i}}$.
  1. 1.
    Offset Approach: In the offset approach, with $Z_{i}$ as the response variable, the weight for contract $i$ is defined as $w_{i}=t_{i}^{2-p}$. Consequently, the MLE for $\zeta$, denoted by $\widehat{\zeta}^{O}$, is:
    $\widehat{\zeta}^{O}=\frac{\sum_{i=1}^{n}t_{i}^{2-p}z_{i}}{\sum_{i=1}^{n}t_{i}^{2-p}}.$ (4.17)
    The sum of estimated premiums is then calculated as:
    $\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{O}$ $=\sum_{i=1}^{n}t_{i}\left(\frac{\sum_{i=1}^{n}t_{i}^{2-p}z_{i}}{\sum_{i=1}^{n}t_{i}^{2-p}}\right)$
    $=\frac{\sum_{i=1}^{n}t_{i}^{2-p}z_{i}}{\sum_{i=1}^{n}t_{i}^{2-p}}\sum_{i=1}^{n}t_{i}.$
    We observe that the estimated premiums sum matches the aggregate observed loss cost only under the conditions that $t_{i}=1$ or $p=1$. Given our assumption that the variance parameter $21<p<$ 2 and $t_{i}<1$, it follows that:
    $\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{O}\neq\sum_{i=1}^{n}t_{i}z_{i}.$
  2. 2.
    Ratio Approach: In the ratio approach, where $Z_{i}$ also serves as the response variable, the weight for contract $i$ is defined as $w_{i}=t_{i}$. The MLE for $\zeta$, denoted by $\widehat{\zeta}^{R}$, is calculated as:
    $\widehat{\zeta}^{R}=\frac{\sum_{i=1}^{n}t_{i}z_{i}}{\sum_{i=1}^{n}t_{i}}.$ (4.18)
    In this case, the aggregate observed loss costs matches the sum of the expected premiums, as shown by:
    $\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{R}$ $=\sum_{i=1}^{n}t_{i}\left(\frac{\sum_{i=1}^{n}t_{i}z_{i}}{\sum_{i=1}^{n}t_{i}}\right)$
    $=\frac{\sum_{i=1}^{n}t_{i}z_{i}}{\sum_{i=1}^{n}t_{i}}\sum_{i=1}^{n}t_{i}$
    $=\sum_{i=1}^{n}t_{i}z_{i}.$
    This result suggests that the ratio approach inherently ensures that the sum of estimated premiums aligns with the aggregate observed loss costs, making it advantageous for achieving balance at the portfolio level.
4.4.2 Heterogeneous Portfolio
We now compare the approaches in the case of a heterogeneous portfolio with explanatory variables. We then incorporate the characteristics vector $\mathbf{X}_{i}$ to model the premium for each contract $i$, allowing for contract-specific premium estimation and accounting for variation across contracts. Thus, the premium in this approach is represented by Equation (4.1). Unlike the homogeneous portfolio, where a closed-form solution for the MLE of $\beta$ was feasible, here the MLE for the parameter vector $\beta$ does not have a closed-form solution due to its dependence on $\mathbf{X}_{i}$. Consequently, we apply the IRLS algorithm (see Section 2.4.1) to obtain the MLE of $\beta$ for both the offset and ratio approaches, as discussed in previous sections.
To better understand the gap between the aggregate observed loss costs and the sum of estimated premiums in each approach, we initialize the IRLS algorithm with values of $\beta$ based on the homogeneous portfolio assumption. This initialization promotes consistency in comparing the offset and ratio approaches and clarifies how incorporating $\mathbf{X}_{i}$ affects the alignment of estimated premiums with aggregate loss costs.
  1. 1.
    Offset Approach: We initialize the IRLS algorithm with:
    $\beta_{0}=\log\left(\widehat{\zeta}^{O}\right),\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \beta_{j}=0,\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ j=1,\ldots,q.$
    By denoting the initial value by $\beta^{O}_{(0)}$, the estimate of the parameter vector at iteration $K$, $\beta^{O}_{(K)}$, is obtained as follows:
    $\beta^{O}_{(K)}=K\beta^{O}_{(0)}+\underbrace{\sum_{k=1}^{K-1}\left(I^{O}_{n}\left(\beta^{O}_{(k)}\right)\right)^{-1}\nabla^{O}\left(\beta^{O}_{(k)}\right)}_{\delta^{O}_{K-1}}=K\beta^{O}_{(0)}+\delta^{O}_{K-1},$
    where $I^{O}_{n}(\cdot)$ is the Fisher information matrix for the offset approach. Therefore, we calculate the estimate premium at iteration $K$, denoted by $\widehat{\zeta}^{O}_{i(K)}$, as follows:
    $\widehat{\zeta}^{O}_{i(K)}=\left(\widehat{\zeta}^{O}\right)^{K}\exp\left\{\mathbf{X}^{T}_{i}\delta^{O}_{K-1}\right\},$
    where $\widehat{\zeta}^{O}$, defined by equation (4.17), represents the annual premium for all insureds in the offset approach for the homogeneous portfolio. We can conclude from this last equation that the sum of the estimated premiums at iteration $K$, denoted as $\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{O}_{i(K)}$, does not necessarily equal the sum of the observed loss costs, $\sum_{i=1}^{n}t_{i}z_{i}$. This discrepancy is typical in the offset approach, given its weighted nature and the way the IRLS updates progress over iterations.
  2. 2.
    Ratio Approach: We initialize the IRLS algorithm with:
    $\beta_{0}=\log\left(\widehat{\zeta}^{R}\right),\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \beta_{j}=0,\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ j=1,\ldots,q.$
    By denoting the initial value by $\beta^{R}_{(0)}$, the estimate of the parameter vector at iteration $K$, $\beta^{R}_{(K)}$, is obtained as follows:
    $\beta^{R}_{(K)}=K\beta^{R}_{(0)}+\underbrace{\sum_{k=1}^{K-1}\left(I^{R}_{n}\left(\beta^{R}_{(k)}\right)\right)^{-1}\nabla^{R}\left(\beta^{R}_{(k)}\right)}_{\delta^{R}_{K-1}}=K\beta^{R}_{(0)}+\delta^{R}_{K-1},$
    where $I^{R}_{n}(\cdot)$ is the Fisher information matrix for the ratio approach. Therefore, we calculate the estimate premium at iteration $K$ as follows:
    $\widehat{\zeta}^{R}_{i(K)}=\left(\widehat{\zeta}^{R}\right)^{K}\exp\left\{\mathbf{X}^{T}_{i}\delta^{R}_{K-1}\right\},$
    where $\widehat{\zeta}^{R}$, defined by equation (4.18), represents the annual premium for all insureds in the ratio approach for the homogeneous portfolio. This allows us to calculate the sum of the estimated premiums at iteration $K$, $t_{i}\widehat{\zeta}^{R}_{i(K)}$, leading to the following equation:
    $\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{R}_{i(K)}$ $=$ $\left(\sum_{i=1}^{n}y_{i}\right)\times\left(\frac{\left(\widehat{\zeta}^{R}\right)^{K-1}}{\sum_{i=1}^{n}t_{i}}\sum_{i=1}^{n}t_{i}\exp\left\{\mathbf{X}^{T}_{i}\delta^{R}_{K-1}\right\}\right)$
    $=$ $\sum_{i=1}^{n}y_{i}\times\epsilon_{(K-1)},$
    highlighting the connection between the total premiums and the aggregate claims $\sum_{i=1}^{n}y_{i}$. That means that unless $\epsilon_{(K-1)}$ equals 1, there is no financial equilibrium for this approach. While the sum of estimated premiums may not exactly match the observed aggregate loss costs, this formulation allows for an analysis of $\epsilon_{(K-1)}$, the gap between these two quantities.
4.4.3 Numerical Illustration
To further compare the observed gaps in both regression approaches, especially for the heterogeneous portfolio where an explicit analytical solution is not available, we conduct a numerical example. The following assumptions are made:
  1. 1.
    Portfolio Setup: In the homogeneous case, we consider $n=100$ independent contracts with no covariates. In the heterogeneous case, we include two risk factors in the mean parameter of each contract, generated from binomial distributions:
    • $\mathbf{x_{1}}\sim\text{Binomial}(100,0.75)$
    • $\mathbf{x_{2}}\sim\text{Binomial}(100,0.15)$
  2. 2.
    Risk Exposure: Risk exposures are randomly drawn from a uniform distribution over the interval $\left[\frac{30}{365},\frac{335}{365}\right]$, and then sorted in increasing order to assign each contract a corresponding rank $i$.
  3. 3.
    Loss Cost Scenarios:
    • Increasing Scenario: Loss costs increase with risk exposure: $y_{i}=i$, for $i=1,\ldots,n$.
    • Decreasing Scenario: Loss costs decrease with risk exposure: $y_{i}=n-i-1$, for $i=1,\ldots,n$.
This small numerical example allows us to highlight some of the analytical results discussed earlier. More importantly, it helps to better understand the behavior of the observed gaps for a heterogeneous portfolio, which may not always be expressed analytically based on the approaches (as discussed in Section 4.4.2). Thus, for the Increasing and the Decreasing scenarios, we define the individual gap for contract $i$ using the following equations:
$\Delta_{i}^{R}=t_{i}\left(z_{i}-\widehat{\zeta}^{R}_{i}\right)\quad\text{and}\quad\Delta_{i}^{O}=t_{i}\left(z_{i}-\widehat{\zeta}^{O}_{i}\right).$
Both $\Delta_{i}^{R}$ and $\Delta_{i}^{O}$ were already used in equation (4.14). For risk exposure in $\left[\frac{30}{365},\frac{335}{365}\right]$, the differences $\Delta_{i}$ are plotted as shown in Figure 2. Depending on the portfolio (homogeneous or heterogeneous), the approach (offset or ratio), and the scenario (increasing or decreasing) being studied, many conclusions can be made from Figure 2.
  • Ratio Approach (with solid and dashed blue lines)
    We begin the analysis by recalling that, as shown in Section 4.4.1, $\sum_{i=1}^{n}\Delta_{i}^{R}=0$ for the homogeneous portfolio. This corresponds to the blue dashed line in both graphs of Figure 2. In comparison, the two curves (solid blue lines) representing the heterogeneous portfolio case are very close to the homogeneous case, indicating that $\sum_{i=1}^{n}\Delta_{i}^{R}$ for the heterogeneous case is expected to be near zero. However, the small jumps observed in both solid blue lines suggest that full financial equilibrium cannot be fully achieved.
  • Offset Approach (with solid and dashed red lines)
    We showed in Section 4.4.1 that the offset approach does not allow for $\sum_{i=1}^{n}\Delta_{i}^{O}=0$, which is exactly what we observe with all the red lines of the homogeneous portfolio. Since the sum of the values of the dashed blue line equals zero, the upward (or downward) shift of the red line indicates that financial equilibrium is not achieved. We also observe the same small jumps for the solid blue line associated with the heterogeneous portfolio. In the scenario where costs increase with exposure, it is interesting to note that the offset approach line is shifted upwards compared to the ratio approach line, while the opposite shift is observed in the decreasing scenario. Thus, we can see that the gap in financial equilibrium for the offset approach will be either positive or negative, depending on whether we are in an increasing or decreasing cost-exposure scenario.
Figure 2: Observed gap by risk exposure (left: Scenario 1, right: Scenario 2)
5 Empirical Illustration
To illustrate the offset and ratio approaches with real insurance data, we utilize a non-random sample from a car insurance database maintained by a major Canadian insurer, spanning 13 consecutive years. The dataset pertains to the province of Ontario and includes over 2 million observations, with each observation representing an annual insurance contract for an individual vehicle. For each contract, the dataset includes key identifiers such as policy number, vehicle ID, and contract start and end dates. The dataset also contains detailed information about the insured individuals and their vehicles, as well as the number and cost of claims associated with each contract. The database is segmented by coverage type, providing insights into third-party liability, collision, and comprehensive claims.
For the purposes of this study, we focus on a single coverage type: collision coverage, which addresses property damage protection for at-fault accidents. To ensure the independence of contracts, we modified the original dataset to include only one vehicle contract per policy. In this revised dataset, each policy corresponds to a unique contract, reducing the dataset to approximately 161,390 contracts.
5.1 Description of Data
5.1.1 Description of Contracts
A standard car insurance contract typically carries a one-year risk exposure; however, cancellations can occur within this period. To offset the administrative costs associated with such cancellations, the insured is subject to a financial penalty for breaching the contract. Several scenarios can lead to mid-term cancellations:
  1. 1.
    The insured sells the vehicle and no longer requires insurance coverage.
  2. 2.
    The insured opts to switch to another insurer, potentially attracted by a lower cost despite the cancellation penalty.
  3. 3.
    The insured experiences a significant loss (e.g., theft or accident) and subsequently changes vehicles, rendering the original insurance coverage unnecessary.
In cases #1 and #3, it could be argued that the situation constitutes a vehicle change rather than an outright cancellation of insurance. In practice, insurers often do not impose penalties for such changes, provided the insured continues coverage with the same insurer for the new vehicle. However, a vehicle change frequently enables the insured to explore options with other insurers, which can quickly lead to scenario #2.
5.1.2 Summary of Data
In Table 1, we present a summary of the data available for the numerical application. It shows that the majority of the observations in the database come from insured individuals covered for a full year. However, a significant portion of the contracts observed in the database corresponds to mid-term cancellations. It can be observed that the average exposure of contracts without cancellations is 1, while contracts with mid-term cancellations have an average exposure of approximately 0.5. The table also indicates the average risk exposure for each group. The last column presents what we refer to as the Loss Cost Reference, which is derived by dividing the average loss cost of each group by the portfolio’s average  [^2] The use of the Loss Cost Reference helps prevent excessive disclosure of confidential information from the insurer who provided us with the data.. A notable difference emerges between the two groups, indicating that insureds with mid-term cancellations tend to have higher insurance damages than those in the other group.
Proportion Average Risk Loss Cost
Group of contracts Exposure Reference
Full exposure 64% 1.00 0.63
Mid-term cancellation 36% 0.51 2.45
Table 1: Descriptive statistics by group of contracts
Figure 3 presents a detailed view of the risk exposure distribution for insured individuals with mid-term cancellations, while the risk exposure for other insured individuals remains constant at 1. The distribution shows that risk exposure is fairly uniform throughout the year; however, there is still considerable variation.
Figure 3: Distribution of risk exposures
To further examine the relationship between risk exposure and loss costs, we grouped contracts by the total number of coverage months, from 1 to 11. For each coverage duration, we calculated the loss cost reference and plotted these values against the average risk exposure, as shown in Figure 4. Insured individuals with full exposure (coverage for 12 months) do not appear in the groups for months 1 to 11; thus, only their average value is displayed in Figure 4 as a dashed line.
Figure 4: Loss cost references by risk exposures
For insured individuals with mid-term cancellations, the loss cost reference curves exhibit a clear downward trend as risk exposure increases, which clearly aligns with the Decreasing Scenario in Section 4.4.3. This seems to imply that policyholders who cancel at the beginning of their contract have more claims, or claims with greater severity, than others. This finding highlights the significant role of risk exposure in explaining contract costs. Although traditional pricing models face challenges in predicting an insured’s risk exposure or the likelihood of cancellation before the end of their contract, incorporating this information into pricing models is crucial for accurately estimating parameters and setting premiums.
5.1.3 Covariates Description
The dataset contains various characteristics for each vehicle and contract. To examine the impact of segmentation on pricing, we selected nine key characteristics as covariates, denoted as $X_{0}$, $X_{1}$, $X_{2}$, $X_{3}$, $X_{4}$, $X_{5}$, $X_{6}$, $X_{7}$, and $X_{8}$ for confidentiality purposes. $X_{0}$ represents a covariate with a constant value of 1 across all contracts, while the other covariates ($X_{1}$ through $X_{8}$) are binary variables. These covariates capture conventional risk factors such as the insured’s gender, age, vehicle usage, and type of vehicle, in line with standard actuarial practices. To simplify the model, we excluded risk factors that are not typically used in pricing models. As such, we believe the resulting pricing model reflects standard industry practices. Descriptive statistics for these covariates are provided in the appendix.
5.2 Offset and Ratio Approaches
To compare the offset and ratio regression approaches, we estimate the parameters using a Tweedie distribution for each segment of the portfolio (insureds with full exposure and those with mid-term cancellation). Furthermore, for each approach (ratio and offset), we calculate the premium for each contract using both methods.
It is worth noting that our dataset is a sample from Boucher and Coulibaly (2024). Accordingly, we applied the same estimated variance parameter, $p=1.42$, as used in their study for premium modelling. As discussed in Section 2, the dispersion parameter does not affect premium modelling; thus, we do not emphasize its role in this analysis.
5.2.1 Estimated Parameters
The estimated parameters $\beta$ for covariates $X_{1}-X_{8}$ can be different for both approaches. To compare the estimated values, we compute a coefficient ratio for each covariate by dividing the estimated coefficient from the offset approach by the corresponding coefficient from the ratio approach. If the estimated values were identical for both approaches, this ratio would be 1 for all covariates. These results are displayed in Figure 5. Without suprise, the estimated coefficients are identical in both the offset and ratio approaches, which was expected because the offset and ratio approaches yield equivalent results when the risk exposure is equal to 1.
On the other hand, for contracts with mid-term cancellations, the estimated coefficients differ noticeably between the offset and ratio approaches. This divergence is anticipated because, as discussed in Section 3, the log-likelihood and gradient functions differ for the two approaches when risk exposure is less than 1. The significant differences, particularly in coefficients associated with covariates $X1$, $X2$, and $X4$, highlight how the choice of approach impacts premium estimation. These differences imply that premiums for certain policyholder profiles will vary substantially based on the chosen approach.
Figure 5: Ratio of estimated parameter vectors
5.2.2 Estimated Premiums
We also compared the distribution of premiums for each approach. To do so, we calculated the premium ratio for each method by dividing the estimated premium for each contract under the offset approach by the premium for the same contract under the ratio approach. The distribution of the ratio of estimated premiums, with box-plots, is presented in Figure 6. For the group of insured individuals covered for a one-year period, we observe identical estimated premiums for all contracts in this group (Figure 6). This outcome is expected, as all contracts in this group have a risk exposure of 1, and, as previously discussed, the offset and ratio approaches yield equivalent results in this situation. In contrast, for contracts with mid-term cancellations, we observe premium ratios greater than 1, indicating that the offset approach produces higher premium estimates compared to the ratio approach. This finding is consistent with the Decreasing Scenario discussed in Section 4.4.3, where the average claim amount decreases as the average risk exposure increases.
Figure 6: Box-plot of the ratio of estimated premiums
5.2.3 Financial Equilibrum Analaysis
Financial equilibrium requires that the aggregate observed loss costs closely align with the sum of estimated premiums at each level of the risk factors in the regression model. To assess this, we calculated the aggregate observed loss costs and the sum of estimated premiums for each level of our eight risk factors, across all contract groups and approaches. These values were then organized in ascending order based on the aggregate observed loss costs, and we computed ratios by dividing the sum of estimated premiums (from both the offset and ratio approaches) by the aggregate observed loss costs. These ratios are plotted against aggregate observed loss costs in Figure 7, with risk classes labeled as integers from 1 to 17.
In the Full exposure group, the observed gap is identical for both the offset and ratio approaches, explained by the identical estimated premiums in these methods. As previously noted, particularly in Section 4.3.1, it is interesting to observe that the sums of predictions and observations by risk class differ even when risk exposure equals one. In the Mid-term cancellation group, however, some differences emerge in the observed gaps between approaches. Here, the ratio approach appears to perform better, with the sum of estimated premiums aligning more closely with the aggregate observed loss costs than in the offset approach.
Figure 7: Observed gaps comparison (left: Full exposure, right: Mid-term cancellation)
5.3 Discussion
Even though neither of the two approaches studied is perfect in every respect, the advantages of using the ratio approach for a Tweedie regression are clear for the insurance portfolio we analyzed. However, we also recognize that the analysis of the approach to be taken has brought to light several important issues, which may lead to different choices or potential solutions depending on the context:
  1. 1.
    Approach Choice and Portfolio Composition:
    As shown in Figure 1, the influence of the chosen approach is relatively minor for portfolios with few mid-term cancellations, such as home insurance or certain commercial insurance lines. In these cases, where most contracts have full-year exposures, the offset and ratio approaches produce similar results. However, for insurance portfolios with a higher proportion of contracts having partial-year exposures — typical in certain types of insurance, such as North American car insurance — the choice between the offset and ratio approaches becomes critical.
  2. 2.
    Impact on Premiums and Policyholder Fairness:
    The different coefficient estimates produced by the offset and ratio approaches imply that premiums for the same insured can vary significantly depending on the approach (see the value of estimated parameters from Figure 5). This difference of premium raises important considerations in a competitive insurance market, as it could influence which policyholder profiles benefit or are disadvantaged by each approach. Understanding how these differences impact policyholder fairness and market competitiveness could help insurers decide which premium structure better aligns with market expectations and policyholder satisfaction.
  3. 3.
    Mid-term Cancellations and Loss Costs:
    Table 1 and Figure 3 demonstrate that mid-term cancellations significantly affect loss costs, highlighting the importance of addressing cancellations more precisely in ratemaking models. Rather than simply incorporating risk exposure within a single pricing model, as done with a Tweedie distribution in this paper, analyzing specific causes of cancellations could enable the development of more tailored approaches. By examining the reasons behind cancellations, insurers could create models that account for different cancellation behaviours and adjust premiums accordingly, potentially leading to more accurate pricing strategies.
6 Conclusion
This study compares the ratio and offset approaches to premium estimation, focusing on how each method incorporates risk exposure. The ratio approach assumes a contract’s premium is directly proportional to its risk exposure, treating the exposure as a weight independent of the response variable’s distribution. In contrast, the offset approach, particularly with a logarithmic link function, incorporates the log of the risk exposure as an offset variable in the premium model. Our findings show that these approaches are not equivalent when the response variable follows a Tweedie distribution.
Through a detailed comparison, we identified the respective strengths and weaknesses of each approach. Additionally, we demonstrated that the offset approach can be reformulated as a weighted regression, structurally similar to the ratio approach. Thus, the distinction between these methods primarily lies in the choice of weights, framing the comparison as one between two weighted regressions.
Our analysis of weight selection revealed that, for any given contract, the weight in the offset approach is consistently greater than that in the ratio approach. This supports our conclusion that the premium estimator from the offset approach generally outperforms the ratio approach, especially when considering bias and variance minimization. However, our assessment of the random gap between aggregate loss costs and the sum of premium estimators suggests that this gap tends to be larger for the offset approach than for the ratio approach.
To evaluate the effectiveness of each approach for pricing, we assessed the balance property, which aligns the sum of estimated premiums with the aggregate actual loss costs. Based on this criterion, the ratio approach performed better than the offset approach. As discussed in Section 5.3, we recommend the ratio approach, particularly in portfolios with diverse risk exposures, due to its ability to better align premiums with actual loss costs. For future research, comparing the premium estimators of both methods using auto-calibration techniques, as suggested by Denuit et al. (2021), Lindholm et al. (2023), Denuit et al. (2024) and Denuit and Trufin (2024), would be valuable, especially given the positive bias observed in both approaches.
Appendix: Descriptive statistics of covariates
Figure 8: Descriptive statistics of all 8 covariates from the database

### v2

@@
 Abstract
-In property and casualty insurance, particularly in automobile insurance, risk exposure is traditionally associated with the coverage duration.
-However, factors such as early contract cancellations demand more precise modelling to ensure accurate premium pricing.
-This study introduces and compares two approaches for modelling total claims (or loss costs) in insurance portfolios with a high proportion of policies that have partial-year exposure: the offset and ratio methods.
-We demonstrate that both approaches can be viewed as weighted regressions under the Tweedie distribution framework.
-Through an analysis based on the financial balance property, we find that the ratio approach outperforms the offset method.
-This comparison is illustrated using an automobile insurance portfolio, where a significant share of policyholders terminate their contracts before the coverage period concludes.
-Keywords Tweedie distribution, weighted regression, offset variable, policy cancellation, stochastic order, positive definite matrix
+In property and casualty insurance, particularly in automobile insurance, risk exposure is commonly assumed to be proportional to the duration of coverage.
+This assumption leads to two standard estimation strategies: the ratio approach, which normalizes the response variable (e.g., claim cost or premium) by the exposure, and the offset approach, which incorporates a transformation of the exposure—typically its logarithm—as a fixed regressor in the mean structure of the model.
+Although both approaches rely on the same proportionality assumption, they are not equivalent when the response variable follows a Tweedie distribution, a framework widely used in insurance analytics.
+In this paper, we show that each approach can be implemented independently and yields a consistent estimator of the true mean parameter vector.
+We then show that the offset approach is asymptotically more efficient than the ratio approach, a result established both theoretically and through simulation studies.
+However, when evaluated from the perspective of portfolio-level financial balance, the ratio approach exhibits superior performance, particularly in the presence of heterogeneous or truncated exposures arising from mid-term policy cancellations.
+These theoretical results are illustrated through an empirical analysis of an automobile insurance portfolio with a high cancellation rate, highlighting the practical implications of model choice for premium estimation under variable exposure conditions.
+Keywords Tweedie distribution, weighted regression, offset variable, policy cancellation, positive definite matrix
 1 Introduction
@@
 (2014)), the premium can be expressed as:
-$\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{T}\beta\right),$ (1.1)
-where $\beta$ is a parameter vector to be estimated, and $\mathbf{X_{i}}=(x_{i,0},\dots,x_{i,q})^{T}$ represents the vector of risk characteristics associated with contract $i$ (with $q$ being a positive integer).
+$\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{\top}\bm{\beta}\right),$ (1.1)
+where $\bm{\beta}$ is a parameter vector to be estimated, and $\mathbf{X_{i}}=(x_{i,0},\dots,x_{i,q})^{\top}$ represents the vector of risk characteristics associated with contract $i$ (with $q$ being a positive integer).
 These risk characteristics—such as age, sex, and the value of the insured goods—are used as covariates in the regression model, allowing the premium to vary based on these factors.
-The function $g(\cdot)$ serves as a link function that connects the score $\mathbf{X}_{i}^{T}\beta$ to the expected value, similar to the approach in Generalized Linear Models (GLMs) (see Nelder and Wedderburn (1972)).
+The function $g(\cdot)$ serves as a link function that connects the score $\mathbf{X}_{i}^{\top}\bm{\beta}$ to the expected value, similar to the approach in Generalized Linear Models (GLMs) (see Nelder and Wedderburn (1972)).
 An analysis of historical motor insurance data reveals significant variability in exposure periods among insured individuals.
@@
     The first approach assumes that the exposure $t_{i}$ is proportional to the mean of the conditional distribution $Y_{i}|\mathbf{X_{i}}$, where $f_{Y_{i}}(\cdot)$ represents the density or probability mass function of $Y_{i}|\mathbf{X_{i}}$, and $y_{i}$ denotes the observed loss cost for contract $i$.
-Under this assumption, the log-likelihood function for estimating the parameter vector $\beta$ is expressed as:
+Under this assumption, the log-likelihood function for estimating the parameter vector $\bm{\beta}$ is expressed as:
     $\ell(\bm{\beta}|y_{1},\ldots,y_{n})=\sum_{i=1}^{n}\log\left(f_{Y_{i}}(y_{i})\right),$
     where the mean of $Y_{i}$ being linearly proportional to $t_{i}$, such that:
-    $\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{T}\beta\right)\times t_{i}.$
+    $\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{\top}\bm{\beta}\right)\times t_{i}.$
     In this framework, $g(\cdot)$ is typically chosen as a logarithmic link function, resulting in:
-    $\exp\left(\mathbf{X}_{i}^{T}\beta\right)t_{i}=\exp\left(\mathbf{X}_{i}^{T}\beta+\log(t_{i})\right),$
+    $\exp\left(\mathbf{X}_{i}^{\top}\bm{\beta}\right)t_{i}=\exp\left(\mathbf{X}_{i}^{\top}\bm{\beta}+\log(t_{i})\right),$
     where $\log(t_{i})$ acts as an offset variable. This method is thus referred to as the offset approach.
@@
     To distinguish it from the offset approach, the mean of $Z_{i}|\mathbf{X_{i}}$ is assumed to no longer depend on $t_{i}$, and the density or probability mass function of $Z_{i}|\mathbf{X_{i}}$ incorporates the risk exposure through this weighting.
-The log-likelihood function for estimating $\beta$ is expressed as:
-    $\ell(\bm{\beta}|z_{1},\ldots,z_{n})=\sum_{i=1}^{n}\log\left(f_{Z_{i}}(z_{i})\right),$
+The log-likelihood function for estimating $\bm{\beta}$ is expressed as:
+    $\ell(\bm{\beta}|z_{1},\ldots,z_{n})=\sum_{i=1}^{n}t_{i}\log\left(f_{Z_{i}}(z_{i})\right),$
     where $f_{Z_{i}}(\cdot)$ is the density or probability mass function of $Z_{i}|\mathbf{X_{i}}$, and the mean is given by:
-    $\mathrm{E}\!\left[Z_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{T}\beta\right),$
+    $\mathrm{E}\!\left[Z_{i}|\mathbf{X_{i}}\right]=g^{-1}\left(\mathbf{X}_{i}^{\top}\bm{\beta}\right),$
     where $g(\cdot)$ a logarithmic link, the expected loss cost for an insured $i$ exposed to risk over a period of $t_{i}$ years is then expressed:
-    $\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=\mathrm{E}\!\left[t_{i}Z_{i}|\mathbf{X_{i}}\right]=\exp\left(\mathbf{X}_{i}^{T}\beta\right)\times t_{i}=\exp\left(\mathbf{X}_{i}^{T}\beta+\log(t_{i})\right),$
+    $\mathrm{E}\!\left[Y_{i}|\mathbf{X_{i}}\right]=\mathrm{E}\!\left[t_{i}Z_{i}|\mathbf{X_{i}}\right]=\exp\left(\mathbf{X}_{i}^{\top}\bm{\beta}\right)\times t_{i}=\exp\left(\mathbf{X}_{i}^{\top}\bm{\beta}+\log(t_{i})\right),$
     which has the same form as the expected loss cost from the offset approach.
-It is important to emphasize that in the offset approach, the random variable $Z_{i}$ can be used to estimate the parameter vector $\beta$, but this method should not be confused with the ratio approach.
+It is important to emphasize that in the offset approach, the random variable $Z_{i}$ can be used to estimate the parameter vector $\bm{\beta}$, but this method should not be confused with the ratio approach.
 Similarly, in the ratio approach, the random variable $Y_{i}$ can also be used, but this method should not ne mistaken for the offset approach.
@@
 Example 1.
-Under the ratio approach, we assume that $Z_{i},i=1,\ldots,n$ follows a Poisson distribution with mean $\zeta_{i}=\exp(\mathbf{X}^{T}_{i}\bm{\beta})$.
-In this framework, it can be shown that the ratio approach corresponds to a Poisson weighted regression and yields the same inference results of the parameter vector $\beta$ as in the offset approach, where $Y_{i},i=1,\ldots,n$ is Poisson distributed with mean $\mu_{i}=t_{i}\exp(\mathbf{X}^{T}_{i}\bm{\beta})=\exp(\mathbf{X}^{T}_{i}\bm{\beta}+\log(t_{i}))$.
+Under the ratio approach, we assume that $Z_{i},i=1,\ldots,n$ follows a Poisson distribution with mean $\zeta_{i}=\exp(\mathbf{X}^{\top}_{i}\bm{\beta})$.
+In this framework, it can be shown that the ratio approach corresponds to a Poisson weighted regression and yields the same inference results of the parameter vector $\bm{\beta}$ as in the offset approach, where $Y_{i},i=1,\ldots,n$ is Poisson distributed with mean $\mu_{i}=t_{i}\exp(\mathbf{X}^{\top}_{i}\bm{\beta})=\exp(\mathbf{X}^{\top}_{i}\bm{\beta}+\log(t_{i}))$.
 Proof.
@@
     $\ell(\bm{\beta}|y_{1},\ldots,y_{n})$ $=\sum_{i=1}^{n}\log\left(\frac{\exp(-\mu_{i})\mu_{i}^{y_{i}}}{y_{i}!}\right)$
-    $=\sum_{i=1}^{n}\left(-t_{i}\exp\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)+y_{i}\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)+y_{i}\log(t_{i})-\log(y_{i}!)\right)$
-    $\propto\sum_{i=1}^{n}\left(-\exp(\mathbf{X}^{T}_{i}\bm{\beta})+z_{i}\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)\right)\times t_{i}.$
+    $=\sum_{i=1}^{n}\left(-t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)+y_{i}\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)+y_{i}\log(t_{i})-\log(y_{i}!)\right)$
+    $\propto\sum_{i=1}^{n}\left(-\exp(\mathbf{X}^{\top}_{i}\bm{\beta})+z_{i}\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)\right)\times t_{i}.$
   • Ratio Approach:
     $\ell(\bm{\beta}|z_{1},\ldots,z_{n})$ $=\sum_{i=1}^{n}t_{i}\log\left(\frac{\exp(-\zeta_{i})\zeta_{i}^{z_{i}}}{z_{i}!}\right)$
-    $=\sum_{i=1}^{n}\left(-t_{i}\exp\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)+t_{i}z_{i}\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)-\log((t_{i}z_{i})!)\right)$
-    $\propto\sum_{i=1}^{n}\left(-\exp(\mathbf{X}^{T}_{i}\bm{\beta})+z_{i}\left(\mathbf{X}^{T}_{i}\bm{\beta}\right)\right)\times t_{i}.$
+    $=\sum_{i=1}^{n}\left(-t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)+t_{i}z_{i}\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)-\log((t_{i}z_{i})!)\right)$
+    $\propto\sum_{i=1}^{n}\left(-\exp(\mathbf{X}^{\top}_{i}\bm{\beta})+z_{i}\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)\right)\times t_{i}.$
 For the Poisson distribution, the offset and the ratio approaches are equivalent, meaning there is no need to choose between them. ∎
 Example 2.
-In the offset approach, we assume that $Y_{i},i=1,\ldots,n$ follows a zero-inflated Poisson distribution characterized by the parameters $\phi$ (for the zero-inflation) and $\mu_{i}=\exp(\mathbf{X}^{T}_{i}\bm{\beta}+\log(t_{i}))$ (from the Poisson distribution).
-In the ratio approach, we assume that $Z_{i},i=1,\ldots,n$ follows a zero-inflated Poisson distribution characterized by the parameters $\phi$ (for the zero-inflation) and $\zeta_{i}=\exp(\mathbf{X}^{T}_{i}\bm{\beta})$ (from the Poisson distribution).
-It can be shown that these two approaches yield different inference results for the estimation of the parameter vector $\beta$.
+In the offset approach, we assume that $Y_{i},i=1,\ldots,n$ follows a zero-inflated Poisson distribution characterized by the parameters $\phi$ (for the zero-inflation) and $\mu_{i}=\exp(\mathbf{X}^{\top}_{i}\bm{\beta}+\log(t_{i}))$ (from the Poisson distribution).
+In the ratio approach, we assume that $Z_{i},i=1,\ldots,n$ follows a zero-inflated Poisson distribution characterized by the parameters $\phi$ (for the zero-inflation) and $\zeta_{i}=\exp(\mathbf{X}^{\top}_{i}\bm{\beta})$ (from the Poisson distribution).
+It can be shown that these two approaches yield different inference results for the estimation of the parameter vector $\bm{\beta}$.
 Proof.
@@
 ∎
-Depending on the distribution used for $Y_{i}$, it is not straightforward to determine whether both approaches are similar.
-Additionally, deciding which approach should be preferred in modelling is not simple.
-Even if the underlying distribution is the same, the parameters obtained from each approach may differ, which could result in varying premiums for each insured.
+Depending on the distribution used for $Y_{i}$, it is not straightforward to determine whether both approaches are equivalent.
+Even when the underlying distribution is assumed to be the same, the estimated parameters obtained from each approach may differ, potentially leading to different premiums for individual policyholders.
+Ideally, if the data-generating process for insurance claims were fully understood, one could determine which modelling approach is most appropriate.
+However, in practice, this is rarely the case, making it difficult to justify one formulation over the other on theoretical grounds alone.
+Consequently, the choice between approaches often relies on a combination of empirical performance and desirable mathematical properties.
+In this context, it is worth noting that Denuit and Trufin (2019) already provide some comparisons between these two approaches.
+In particular, they show that the offset approach can be interpreted as a weighted regression when the response variable follows a Tweedie distribution.
+Building on their work, we extend the comparison further by emphasizing the role of stochastic orders and the principle of financial balance, which provide additional insights into the structural differences between the two modeling approaches.
 1.2 Structure of the paper
-This paper emphasizes the differences between the offset and ratio regression approaches in modelling contract loss costs using the Tweedie distribution, rather than focusing on the number of claims.
-Section 2 introduces the notation and model assumptions, detailing the key properties of the Tweedie distribution and the parameter inference methods employed.
-Section 3 explains the offset and ratio approaches, as well as the link between them.
+This paper emphasizes the differences between the offset and ratio regression approaches in modeling contract loss costs using the Tweedie distribution, rather than focusing on claim frequency.
+Section 2 introduces the notation and model assumptions, detailing key properties of the Tweedie distribution and the parameter inference methods used.
+It also provides the technical framework required to establish the equivalence results and the asymptotic comparisons developed in the remainder of the paper.
+Section 3 then explains the offset and ratio approaches, as well as the link between them.
 We show that both approaches can be viewed as weighted regressions, where only the form of the weight differentiates the approaches.
-An analysis of asymptotic expectations, stochastic dominance, and the gap between aggregate loss costs and estimated premiums using simulated data is presented in Section 4.
-In Section 5, an empirical illustration of both approaches is provided, based on a sample of automobile insurance data from a Canadian insurer.
-The last section concludes.
+Section 4 presents an analysis of the consistency of the parameter vector estimators for both approaches, demonstrates the asymptotic efficiency of the offset-approach parameter vector estimator, and examines the gap between observed aggregate loss costs and the sum of estimated premiums based on simulated data.
+In Section 5, we provide an empirical illustration of both approaches using a sample of automobile insurance data from a Canadian insurer.
+Finally, Section 6 concludes the paper.
 2 GLM with Tweedie-distributed Response Variable
 The loss cost of a contract is the total amount of claims, calculated as the aggregate sum of individual claim costs.
-Formally, let $N$ denote the he total number of claims for a contract, and let $Z_{k}$ represent the cost of the $k$-th claim, where $k,k=1,\ldots,N$ if $N>0$.
+Formally, let $N$ denote the total number of claims for a contract, and let $Z_{k}$ represent the cost of the $k$-th claim, where $k,k=1,\ldots,N$ if $N>0$.
 Thus, assuming $Y$ denotes the loss cost of a contract, as $Y$ is obtained as follows:
@@
 As a result, insurers often use the Tweedie distribution to model a contract premium when the loss cost is the response variable.
-The rationale for using the Tweedie distribution is its ability to represent a mixture distribution (both continuous and discrete), as noted by Tweedie et al.
-(1984).
+The rationale for using the Tweedie distribution is its ability to represent a mixture distribution (both continuous and discrete), as noted by Tweedie and others (1984).
 Additionally, Delong et al.
 (2021) provides a comprehensive summary of this mixture distribution, with relevant applications in actuarial science.
-Accordingly, we consider the Tweedie distribution as a mixture distribution to model the premium using the loss cost as the response variable.
-The justification for employing the GLM theory in this context is based on the fact that the Tweedie distribution, as a mixture distribution, belongs to the linear exponential family of distributions, as discussed in Tweedie et al.
-(1984) and Jørgensen (1987).
-Before presenting the GLM approach in the context of the Tweedie distribution, we review the Tweedie distribution by referring to the works of Tweedie et al.
-(1984), Jørgensen (1987), and Delong et al.
+Accordingly, we model the premium using the loss cost as the response variable, assuming that the latter follows a Tweedie distribution.
+This distribution is viewed as a compound Poisson–Gamma mixture and belongs to the linear exponential family, which justifies the use of the Generalized Linear Model (GLM) framework (see Tweedie, 1984; Jørgensen, 1987).
+Before introducing the GLM approach with the Tweedie distribution, we briefly review its properties by referring to Tweedie and others (1984), Jørgensen (1987), and Delong et al.
 (2021).
+It is also important to note that several results presented here can also be found in Wüthrich and Merz (2023) and Denuit and Trufin (2019); however, this review is essential for a proper understanding of the offset and ratio regression approaches.
 2.1 Exponential Family Form of Tweedie Distribution
@@
   • $a(\cdot)$ denotes a function that does not depend on $\mu$.
-It is important to note that assuming the variance parameter lies within the interval ]1,2[]1,2[] 1 , 2 [ ensures that the Tweedie distribution density is defined as described above.
-As noted by Tweedie et al.
-(1984) and Jørgensen (1987), the Tweedie distribution can be generalized for any positive variance parameter $p>0$.
-However, they also mention that when the variance parameter lies within the interval ]0,1[]0,1[] 0 , 1 [, the Tweedie distribution does not belong to the exponential family.
+It is important to note that assuming the variance parameter lies within the interval ]1,2[]1,2[ ensures that the Tweedie distribution density is defined as described above.
+As noted by Tweedie and others (1984) and Jørgensen (1987), the Tweedie distribution can be generalized for any positive variance parameter $p>0$.
+However, they also mention that when the variance parameter lies within the interval ]0,1[]0,1[, the Tweedie distribution does not belong to the exponential family.
 The density function $f(\cdot)$ mentioned above is expressed in the form of the exponential family density function.
@@
 It should also be noted that $f(\cdot)$ lacks a specific form because the function $a(\cdot)$ is unknown.
-To determine $a(.)$, one can refer to Delong et al.
+To determine $a(\cdot)$, one can refer to Delong et al.
 (2021), which provides a comprehensive overview of the Tweedie distribution.
@@
 This paper will analyze this gap in detail.
-For the remainder of this paper, we denote the parameters of the Tweedie distribution by $(\mu,w,\phi,p)$ and we define $\mathcal{T}$ as the set of Tweedie distributions $(\mu,w,\phi,p)$ where the variance parameter $p$ lies within the interval ]1,2[]1,2[] 1 , 2 [.
+For the remainder of this paper, we denote the parameters of the Tweedie distribution by $(\mu,w,\phi,p)$ and we define $\mathcal{T}$ as the set of Tweedie distributions $(\mu,w,\phi,p)$ where the variance parameter $p$ lies within the interval ]1,2[]1,2[.
 2.2 Notations
@@
   • $x_{i,j}$: The observed characteristic for contract $i$ corresponding to risk factor $\bm{x_{j}}$, where $j=1,\ldots,q$.
-  • $\mathbf{X}_{i}$: The vector of risk characteristics associated with contract $i$, defined as $\mathbf{X}_{i}=\left(1,x_{i,1},\ldots,x_{i,q}\right)^{T}$.
+  • $\mathbf{X}_{i}$: The vector of risk characteristics associated with contract $i$, defined as $\mathbf{X}_{i}=\left(x_{i,0},x_{i,1},\ldots,x_{i,q}\right)^{\top}$, where $x_{i,0}=1,\forall i=1,\ldots,n$.
 Finally, the matrix $\bm{X}$, representing the risk factors, is given by:
@@
   2. 2.
-    $Y_{i_{1}}|\mathbf{X}_{i_{1}}\perp\!\!\!\perp Y_{i_{2}}|\mathbf{X}_{i_{2}},\forall i_{1},i_{2}=1,\ldots,n\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ i_{1}\neq i_{2}$.
+    $Y_{i_{1}}|\mathbf{X}_{i_{1}}\perp\!\!\!\perp Y_{i_{2}}|\mathbf{X}_{i_{2}},\forall i_{1},i_{2}=1,\ldots,n\penalty 10000\ \text{such that}\penalty 10000\ i_{1}\neq i_{2}$.
   3. 3.
@@
 The third assumption is essential for calculating the variance of the estimator of $\mu_{i}$.
-The concept of GLM revolves around statistical inference on $\mu_{i}$, which is considered the premium for contract $i$ premium in insurance pricing.
+The concept of GLM revolves around statistical inference on $\mu_{i}$, which is considered the premium for contract $i$ in insurance pricing.
 This consideration is justified by the fact that the mean parameter $\mu_{i}$, regardless of the distribution of $Y_{i}|\mathbf{X}_{i}$, corresponds to the constant closest to $Y_{i}|\mathbf{X}_{i}$ in terms of Euclidean distance (for further details, see Denuit et al.
@@
 Thus, for modelling the premium of contract $i$, we assume that the logarithm of this premium ($\mu_{i}$) linearly depends on the characteristics vector $\mathbf{X}_{i}$ as follows:
-$\mu_{i}$ $=\exp\left(\bm{X}^{T}_{i}\beta\right)=\exp\left(\beta_{0}\right)\prod_{j=1}^{q}\exp\left(\beta_{j}x_{i,j}\right)=b_{0}\times\prod_{j=1}^{q}r_{j},$ (2.2)
+$\mu_{i}$ $=\exp\left(\bm{X}^{\top}_{i}\bm{\beta}\right)=\exp\left(\beta_{0}\right)\prod_{j=1}^{q}\exp\left(\beta_{j}x_{i,j}\right)=b_{0}\times\prod_{j=1}^{q}r_{j},$ (2.2)
 where:
-  • $\beta$: The parameter vector, defined as $\beta=(\beta_{0},\beta_{1},\ldots,\beta_{q})^{T}$ where $\beta_{j}\in\mathbf{R}$ for $j=0,1,\ldots,q$;
+  • $\bm{\beta}$: The parameter vector, defined as $\bm{\beta}=(\beta_{0},\beta_{1},\ldots,\beta_{q})^{\top}$ where $\beta_{j}\in\mathbf{R}$ for $j=0,1,\ldots,q$;
   • $b_{0}$: The basic premium, defined as $b_{0}=\exp\left(\beta_{0}\right)$.
@@
 While the basic premium $b_{0}$ is consistent across all contracts, the relativity factors depend on the specific risk characteristics of each contract.
-Consequently, statistical inference on $\mu_{i}$ is equivalent to estimating the parameter vector $\beta$.
+Consequently, statistical inference on $\mu_{i}$ is equivalent to estimating the parameter vector $\bm{\beta}$.
 To conclude, it is pertinent to introduce an interesting property of the Tweedie distribution, as discussed in Denuit and Trufin (2019), which will be valuable for further developments.
 Property 2.3.1 (Tweedie Invariance Property).
-If a random variable $Z$ is Tweedie-distributed with parameters $(\mu,w,\phi,p)$ and belongs to the set $\mathcal{T}$, then $tZ$, for any positive $t$, is also Tweedie-distributed with parameters $(t\mu,\frac{w}{t^{2-p}},\phi,p)$ and belongs to the same set $\mathcal{T}$.
-Conversely, if $tZ$, for any positive $t$, is Tweedie distributed with parameters $(\mu,w,\phi,p)$ and belongs to the set $\mathcal{T}$, then $Z$ is also Tweedie-distributed with parameters$(\frac{\mu}{t},wt^{2-p},\phi,p)$ and belongs to the same set $\mathcal{T}$.
-Formally, we have:
+[…]
 $Z\in\mathcal{T}\Longleftrightarrow tZ\in\mathcal{T},\forall t>0.$
 2.4 Log-likelihood Function
-The GLM employs the maximum likelihood approach to estimate the parameter vector $\beta$ (Nelder and Wedderburn, 1972).
-The maximum likelihood approach maximizes the likelihood function, which depends on the parameter vector $\beta$ given the observations $y_{1},\ldots,y_{n}$.
+The GLM employs the maximum likelihood approach to estimate the parameter vector $\bm{\beta}$ (Nelder and Wedderburn, 1972).
+The maximum likelihood approach maximizes the likelihood function, which depends on the parameter vector $\bm{\beta}$ given the observations $y_{1},\ldots,y_{n}$.
 For a detailed discussion on the theory of maximum likelihood, see Casella and Berger (2024).
@@
 For the Tweedie case, the gradient function is derived as follows:
-$\nabla(\beta)$ $=\left(\frac{\partial\ell(\beta_{0},\ldots,\beta_{q})}{\partial\mu_{i}}\frac{\partial\mu_{i}}{\partial\beta_{0}},\ldots,\frac{\partial\ell(\beta_{0},\ldots,\beta_{q})}{\partial\mu_{i}}\frac{\partial\mu_{i}}{\partial\beta_{j}}\right)^{t}$
-$=\frac{1}{\phi}\left(\sum_{i=1}^{n}w_{i}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}},\ldots,\sum_{i=1}^{n}w_{i}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}}x_{i,q}\right)^{t}$
-$=\frac{1}{\phi}\bm{X}^{t}\bm{D}\bm{R};$
+$\nabla(\bm{\beta})$ $=\left(\frac{\partial\ell(\beta_{0},\ldots,\beta_{q})}{\partial\mu_{i}}\frac{\partial\mu_{i}}{\partial\beta_{0}},\ldots,\frac{\partial\ell(\beta_{0},\ldots,\beta_{q})}{\partial\mu_{i}}\frac{\partial\mu_{i}}{\partial\beta_{j}}\right)^{T}$
+$=\frac{1}{\phi}\left(\sum_{i=1}^{n}w_{i}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}},\ldots,\sum_{i=1}^{n}w_{i}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}}x_{i,q}\right)^{\top}$
+$=\frac{1}{\phi}\bm{X}^{\top}\bm{D}\bm{R};$
 The vector $\bm{R}$ and the matrix $\bm{D}$ are defined as follows:
@@
 We also recall the definition of positive definite matrices and their properties, as discussed by Horn and Johnson (2012).
-Let $\beta_{(k)}$ denote the estimate of $\beta$ at iteration $k$, The IRLS algorithm is defined as follows:
-$\beta_{(k+1)}=\beta_{(k)}+\left(I_{n}\left(\beta_{(k)}\right)\right)^{-1}\nabla\left({\beta_{(k)}}\right);$ (2.6)
+Let $\bm{\beta}_{(k)}$ denote the estimate of $\bm{\beta}$ at iteration $k$, The IRLS algorithm is defined as follows:
+$\bm{\beta}_{(k+1)}=\bm{\beta}_{(k)}+\left(I_{n}\left(\bm{\beta}_{(k)}\right)\right)^{-1}\nabla\left({\bm{\beta}_{(k)}}\right);$ (2.6)
 $I_{n}(\cdot)$ is the Fisher information matrix and is defined as follows:
-$I_{n}(\beta)$ $=-\mathrm{E}\!\left[\text{Hess}\left({\beta}\right)\right]$
+$I_{n}(\bm{\beta})$ $=-\mathrm{E}\!\left[\text{Hess}\left({\bm{\beta}}\right)\right]$
 $=\frac{1}{\phi}\left(\sum_{i=1}^{n}w_{i}\mu^{2-p}_{i}x_{i,j_{1}}x_{i,j_{2}}\right)_{{}_{j_{1},j_{2}=0,\ldots,q}}$
-$=\frac{1}{\phi}\bm{X}^{t}\bm{D}\bm{X},$
-where $\text{Hess}\left({\beta}\right)$ is the Hessian matrix.
+$=\frac{1}{\phi}\bm{X}^{\top}\bm{D}\bm{X},$
+where $\text{Hess}\left({\bm{\beta}}\right)$ is the Hessian matrix.
 Referring to equation (2.6), we note that the IRLS algorithm relies on the invertibility of the Fisher Information Matrix.
@@
 A matrix $M\in\mathcal{M}_{m,m}$ is positive definite ($M\succ 0$), if the following condition holds:
-$x^{t}Mx>0;\forall x\in\mathbb{R}^{m}.$
+$x^{\top}Mx>0;\forall x\in\mathbb{R}^{m}.$
 Property 2.4.1.
@@
 $\forall A,B\in\mathcal{P}_{m}:A-B\succ 0$ $\Longleftrightarrow B^{-1}-A^{-1}\succ 0;$ (2.8)
-$\forall A\in\mathcal{P}_{m},\forall X\in\mathcal{M}_{m,q}:X^{t}AX\succ 0$ $\Longleftrightarrow X\leavevmode\nobreak\ \text{has rank}\leavevmode\nobreak\ q.$ (2.9)
+$\forall A\in\mathcal{P}_{m},\forall X\in\mathcal{M}_{m,q}:X^{\top}AX\succ 0$ $\Longleftrightarrow X\penalty 10000\ \text{has rank}\penalty 10000\ q.$ (2.9)
 Property (2.7) serves as a necessary and sufficient condition for proving the invertibility of a matrix.
@@
 Finally, if the weight of each contract is known and the variance parameter is also known, the GLM in the Tweedie case involves using the IRLS algorithm to obtain an estimate of the parameters vector.
-$\left(I_{n}(\beta)\right)^{-1}\nabla\left(\beta\right)=\left(X^{t}DX\right)^{-1}X^{t}DR.$ (2.10)
+$\left(I_{n}(\bm{\beta})\right)^{-1}\nabla\left(\bm{\beta}\right)=\left(X^{\top}DX\right)^{-1}X^{\top}DR.$ (2.10)
 It is also noteworthy that the IRLS algorithm is implemented in the R software through the glm function.
@@
 3.1 Offset Approach
-As mentioned in Section 1.1, the statistical inference on the parameter vector $\beta$ in offset approach is performed with the lost costs $Y_{1},\ldots,Y_{n}$.
-To account for the logarithm of a contract’s risk exposure as the offset variable, we assume that $Y_{i}|\mathbf{X}_{i}$ follows the Tweedie distribution with parameters $(\mu_{i},1,\phi,p)$ where $\mu_{i}$ is defined as $\mu_{i}=t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right)$.
+As mentioned in Section 1.1, the statistical inference on the parameter vector $\bm{\beta}$ in offset approach is performed with the loss costs $Y_{1},\ldots,Y_{n}$.
+To account for the logarithm of a contract’s risk exposure as the offset variable, we assume that $Y_{i}|\mathbf{X}_{i}$ follows the Tweedie distribution with parameters $(\mu_{i},1,\phi,p)$ where $\mu_{i}$ is defined as $\mu_{i}=t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right)$.
 We also denote an estimate of $\mu_{i}$ in the offset regression by $\widehat{\mu}^{O}_{i}$, obtained as follows:
-$\widehat{\mu}^{O}_{i}=\exp\left(\mathbf{X}^{T}_{i}\widehat{\beta}^{O}+\log(t_{i})\right),$ (3.1)
-where $\widehat{\beta}^{O}$, an estimate of $\beta$ in offset approach, is obtained using the IRLS algorithm with the following quantities:
+$\widehat{\mu}^{O}_{i}=\exp\left(\mathbf{X}^{\top}_{i}\widehat{\bm{\beta}}^{O}+\log(t_{i})\right),$ (3.1)
+where $\widehat{\bm{\beta}}^{O}$, an estimate of $\bm{\beta}$ in offset approach, is obtained using the IRLS algorithm with the following quantities:
   • Log-likelihood Function:
@@
   • Gradient Function:
-    $\nabla^{O}(\beta)$ $=\frac{1}{\phi}\left(\sum_{i=1}^{n}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}},\ldots,\sum_{i=1}^{n}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}}x_{i,q}\right)^{t}$ (3.3)
-    $=\frac{1}{\phi}\bm{X}^{t}\bm{D}^{O}\bm{R}^{O};$ (3.4)
+    $\nabla^{O}(\bm{\beta})$ $=\frac{1}{\phi}\left(\sum_{i=1}^{n}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}},\ldots,\sum_{i=1}^{n}\frac{y_{i}-\mu_{i}}{\mu^{p-1}_{i}}x_{i,q}\right)^{\top}$ (3.3)
+    $=\frac{1}{\phi}\bm{X}^{\top}\bm{D}^{O}\bm{R}^{O};$ (3.4)
     The vector $\bm{R}^{O}$ and the matrix $\bm{D}^{O}$ are defined as follows:
-    $\bm{D}^{O}=\operatorname{diag}\left(t^{2-p}_{i}\exp\left((2-p)\mathbf{X}^{T}_{i}\beta\right)\right)_{i=1,...,n},\bm{R}^{O}=\left(\frac{y_{i}}{\mu_{i}}-1\right)_{i=1,...,n}.$
+    $\bm{D}^{O}=\operatorname{diag}\left(t^{2-p}_{i}\exp\left((2-p)\mathbf{X}^{\top}_{i}\bm{\beta}\right)\right)_{i=1,...,n},\bm{R}^{O}=\left(\frac{y_{i}}{\mu_{i}}-1\right)_{i=1,...,n}.$
 3.2 Ratio Approach
@@
 To account for the risk exposure $t_{i}$ as the weight of contract $i$, we also assume $Z_{i}|\mathbf{X}_{i}$ follows a Tweedie distribution with parameters $(\zeta_{i},t_{i},\phi,p)$, where the mean parameter $\zeta_{i}$ is assumed to be:
-$\zeta_{i}=\exp\left(\mathbf{X}^{T}_{i}\beta\right),$ (3.5)
-and $\widehat{\zeta_{i}}$ is an estimate of $\zeta_{i}$, equal to $\widehat{\zeta_{i}}=\exp\left(\mathbf{X}^{T}_{i}\widehat{\beta}^{R}\right)$.
-The vector $\widehat{\beta}^{R}$, an estimate of $\beta$ in ratio approach, is obtained using the IRLS algorithm with the following quantities:
+$\zeta_{i}=\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),$ (3.5)
+and $\widehat{\zeta_{i}}$ is an estimate of $\zeta_{i}$, equal to $\widehat{\zeta_{i}}=\exp\left(\mathbf{X}^{\top}_{i}\widehat{\bm{\beta}}^{R}\right)$.
+The vector $\widehat{\bm{\beta}}^{R}$, an estimate of $\bm{\beta}$ in ratio approach, is obtained using the IRLS algorithm with the following quantities:
   • Log-likelihood Function: This function is constructed with the observations $z_{1},\ldots,z_{n}$ where $\zeta_{i}=\frac{y_{i}}{t_{i}},i=1,\ldots,n$, as follows:
@@
   • Gradient Function:
-    $\nabla^{R}(\beta)$ $=\frac{1}{\phi}\left(\sum_{i=1}^{n}t_{i}\frac{z_{i}-\zeta_{i}}{\zeta^{p-1}_{i}},\ldots,\sum_{i=1}^{n}t_{i}\frac{z_{i}-\zeta_{i}}{\zeta^{p-1}_{i}}x_{i,q}\right)^{t}$ (3.7)
-    $=\frac{1}{\phi}\bm{X}^{t}\bm{D}^{R}\bm{R}^{R};$ (3.8)
+    $\nabla^{R}(\bm{\beta})$ $=\frac{1}{\phi}\left(\sum_{i=1}^{n}t_{i}\frac{z_{i}-\zeta_{i}}{\zeta^{p-1}_{i}},\ldots,\sum_{i=1}^{n}t_{i}\frac{z_{i}-\zeta_{i}}{\zeta^{p-1}_{i}}x_{i,q}\right)^{\top}$ (3.7)
+    $=\frac{1}{\phi}\bm{X}^{\top}\bm{D}^{R}\bm{R}^{R};$ (3.8)
     The vector $\bm{R}^{R}$ and the matrix $\bm{D}^{R}$ are defined as follows:
-    $\bm{D}^{R}=\operatorname{diag}\left(t_{i}\exp\left\{(2-p)\mathbf{X}^{T}_{i}\beta\right\}\right)_{i=1,...,n},\bm{R}^{R}=\left(\frac{z_{i}}{\zeta_{i}}-1\right)=\left(\frac{y_{i}}{\mu_{i}}-1\right)_{i=1,...,n},$
-    where $\mu_{i}=t_{i}\zeta_{i}=t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right).$
+    $\bm{D}^{R}=\operatorname{diag}\left(t_{i}\exp\left\{(2-p)\mathbf{X}^{\top}_{i}\bm{\beta}\right\}\right)_{i=1,...,n},\bm{R}^{R}=\left(\frac{z_{i}}{\zeta_{i}}-1\right)=\left(\frac{y_{i}}{\mu_{i}}-1\right)_{i=1,...,n},$
+    where $\mu_{i}=t_{i}\zeta_{i}=t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right).$
 3.3 Link between Offset and Ratio Approaches
@@
     The offset approach viewed in the form of the ratio approach
-    In the offset approach, $Y_{i}|\mathbf{X}^{T}_{i}$ follows a Tweedie distribution with parameters $(t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)$.
+    In the offset approach, $Y_{i}|\mathbf{X}^{\top}_{i}$ follows a Tweedie distribution with parameters $(t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),1,\phi,p)$.
 We can examine the ratio of the loss cost to the risk exposure $t_{i}$, denoted as $Z_{i}=\frac{Y_{i}}{t_{i}}$, and apply the Tweedie Invariance Property.
-It can then be demonstrated that $Z_{i}|\mathbf{X}^{T}_{i}$ follows a Tweedie distribution with parameters $(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{2-p},\phi,p)$.
+It can then be demonstrated that $Z_{i}|\mathbf{X}^{\top}_{i}$ follows a Tweedie distribution with parameters $(\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i}^{2-p},\phi,p)$.
 More formally, we have the following equivalence:
-    $Y|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)\equiv Z|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{2-p},\phi,p).$
+    $Y|\mathbf{X}^{\top}_{i}\sim\text{Tweedie}(t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),1,\phi,p)\equiv Z|\mathbf{X}^{\top}_{i}\sim\text{Tweedie}(\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i}^{2-p},\phi,p).$
     Using equation (2.5) for the calculation of the matrix $\bm{D}$, we can demonstrate that the two Tweedie distributions introduced earlier share the same expression:
-    $\bm{D}=\operatorname{diag}\left(t^{2-p}_{i}\exp\left((2-p)\mathbf{X}^{T}_{i}\beta\right)\right)_{i=1,\ldots,n}\equiv\bm{D}^{O}$
-    This result provides an additional verification that these two regression approaches are equivalent, leading to identical inference outcomes for the parameter vector $\beta$.
+    D = diag( t^2 - p_i exp( (2 - p) X^⊤_i β ) )_i= 1,…,n ≡D^O
+    This result provides an additional verification that these two regression approaches are equivalent, leading to identical inference outcomes for the parameter vector $\bm{\beta}$.
   2. 2.
     The ratio approach viewed in the form of the offset approach
-    Similarly, in the ratio approach, where $Z_{i}|\mathbf{X}^{T}_{i}$ is assumed to follow a Tweedie distribution with parameters $(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)$, we can apply the Tweedie Invariance Property to demonstrate that $Y_{i}|\mathbf{X}^{T}_{i}$ follows a Tweedie distribution with parameters $(t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{p-1},\phi,p)$.
+    Similarly, in the ratio approach, where $Z_{i}|\mathbf{X}^{\top}_{i}$ is assumed to follow a Tweedie distribution with parameters $(\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i},\phi,p)$, we can apply the Tweedie Invariance Property to demonstrate that $Y_{i}|\mathbf{X}^{\top}_{i}$ follows a Tweedie distribution with parameters $(t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i}^{p-1},\phi,p)$.
 This equivalence is expressed as:
-    $Z|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)\equiv Y|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(t_{i}\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{p-1},\phi,p).$
+    $Z|\mathbf{X}^{\top}_{i}\sim\text{Tweedie}(\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i},\phi,p)\equiv Y|\mathbf{X}^{\top}_{i}\sim\text{Tweedie}(t_{i}\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i}^{p-1},\phi,p).$
     For each form of the Tweedie distributions, we also have the following equivalence for the matrix $\bm{D}$:
-    $\bm{D}=\operatorname{diag}\left(t_{i}\exp\left((2-p)\mathbf{X}^{T}_{i}\beta\right)\right)_{i=1,\ldots,n}\equiv\bm{D}^{R}.$
+    $\bm{D}=\operatorname{diag}\left(t_{i}\exp\left((2-p)\mathbf{X}^{\top}_{i}\bm{\beta}\right)\right)_{i=1,\ldots,n}\equiv\bm{D}^{R}.$
 The previous results facilitate a direct comparison of the two approaches. By focusing on the random variable $Z$, the choice is between the following two distributions:
-$\text{Tweedie}(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i}^{2-p},\phi,p)\ \ \ \text{{versus}}\ \ \ \text{Tweedie}(\exp\left(\mathbf{X}^{T}_{i}\beta\right),t_{i},\phi,p)$
+$\text{Tweedie}(\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i}^{2-p},\phi,p)\ \ \ \text{{versus}}\ \ \ \text{Tweedie}(\exp\left(\mathbf{X}^{\top}_{i}\bm{\beta}\right),t_{i},\phi,p)$
 We observe that the choice of approach depends solely on the form of the weights $w_{i}$, as all other parameters remain unchanged.
@@
 This equivalence between the methods is also evident in the equality between $\bm{D}^{R}$ and $\bm{D}^{O}$ when $t=1$:
-$\bm{D}^{R}=\bm{D}^{O}=\operatorname{diag}\left(\exp\left((2-p)\mathbf{X}^{T}_{i}\beta\right)\right)_{i=1,...,n}.$
+$\bm{D}^{R}=\bm{D}^{O}=\operatorname{diag}\left(\exp\left((2-p)\mathbf{X}^{\top}_{i}\bm{\beta}\right)\right)_{i=1,...,n}.$
 We can also interpret Figure 1 by varying the variance parameter $p$.
@@
 4 Comparison of the Approaches
-According to our analysis presented in Section 3.3, for a given value of the variance parameter $p$, the difference between weights $w_{i}^{O}$ and $w_{i}^{R}$, of the offset and ratio approaches respectively, depends on the risk exposure $t_{i}$.
-This difference is significant for contracts with shorter exposure periods, while it becomes negligible for those with exposure periods close to one year.
-However, a limitation of this analysis is that it does not conclusively determine which approach may be more appropriate for insurance pricing.
-Our strategy to compare approaches involves utilizing the distribution of the maximum likelihood estimators of the parameter vectors in the offset and ratio approaches to compare the premium estimators [^1] We will use the expression premium estimator instead of estimated premium throughout this discussion, as the estimated premium refers to the value of a single observation derived from the premium estimator, which is a more general mathematical expression.
-in these two regression methods.
-We begin by comparing the asymptotic expected values and variances of the premium estimators in both regression approaches.
-This comparison utilizes the positive definite matrix property described in equation (2.8), which enables a direct evaluation of the asymptotic covariance matrices of the parameter vector estimators in the offset and ratio regression approaches.
-We conclude our analysis by applying the concept of stochastic order (or stochastic dominance), as discussed in Shaked and Shanthikumar (2007).
-4.1 Notations and assumptions
-We continue to consider $n$ independent contracts, using the same notations and assumptions as in previous sections.
-To facilitate comparison, we define the response variable consistently across both approaches:
-  • Offset approach: $Z|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(\zeta_{i},t^{2-p}_{i},\phi,p)$ for $i=1,\ldots,n$,
-  • Ratio approach: $Z|\mathbf{X}^{T}_{i}\sim\text{Tweedie}(\zeta_{i},t_{i},\phi,p)$ for $i=1,\ldots,n$,
-where $\zeta_{i}$ is the premium for the $i$-th contract without adjustment for risk exposure and is linked to the score $\mathbf{X}^{T}_{i}\beta$ as specified in Equation (4.1):
-$\zeta_{i}=\exp\left(\mathbf{X}^{T}_{i}\beta\right).$ (4.1)
-Using this formulation, the premium estimators for $\zeta_{i}$ in the offset and ratio approaches, denoted by $\bm{\widehat{\zeta}^{O}}_{i}$ and $\bm{\widehat{\zeta}^{R}}_{i}$, are obtained as follows:
-$\bm{\widehat{\zeta}^{O}}_{i}$ $=\exp\left(\mathbf{X}^{T}_{i}\bm{\widehat{\beta}^{O}}\right);$ (4.2)
-$\bm{\widehat{\zeta}^{R}}_{i}$ $=\exp\left(\mathbf{X}^{T}_{i}\bm{\widehat{\beta}^{R}}\right);$ (4.3)
-Where $\bm{\widehat{\beta}^{O}}$ and $\bm{\widehat{\beta}^{R}}$ are the estimators of $\beta$ in the offset and ratio approaches, respectively.
-4.2 Expected Values and Variances Comparison of Premium Estimators
-To compare the offset and ratio regression approaches, a valuable direction is to examine the entire distribution of the premium estimator for each approach.
-However, comparing the expected values and variances of the premium estimators may be sufficient for this analysis.
-Indeed, when evaluating competing estimators, the optimal choice is generally the one with the smallest bias and variance.
-Therefore, our quantities of interest are the expected values and variances of the premium estimators for both approaches, as outlined in the following propositions.
+According to the analysis presented in Section 3.3, for a fixed value of the variance parameter $p$, the difference between the weights $w_{i}^{O}$ and $w_{i}^{R}$ associated with the offset and ratio approaches, respectively, depends on the risk exposure $t_{i}$.
+This difference is substantial for contracts with short exposure periods, while it becomes negligible for contracts with exposure periods close to one year.
+However, a limitation of this analysis is that it does not, by itself, allow us to determine which approach is more appropriate for insurance pricing.
+To address this issue, we proceed in two steps.
+First, we compare the parameter vector estimators derived under the two approaches, relying on the quasi-maximum likelihood framework of White (1982).
+Second, using a numerical illustration, we highlight the practical merits of the ratio approach relative to the offset approach in the context of automobile insurance pricing.
+To this end, we consider as the response variable the loss cost $Y$ per unit of risk exposure $t$, namely $Z=\frac{Y}{t}$, and assume that the true data-generating process of $Z$ is characterized by a density function $g$.
+For all $\bm{x}\in\mathbb{R}^{q+1}$ and $t\in(0,1]$, representing the covariate vector (risk characteristics) and the risk exposure, respectively, the true expected value of $Z$ is given by
+$\zeta^{\text{True}}=\int_{0}^{+\infty}z\,g\!\left(z\mid\bm{\beta}^{\text{True}},t\right)\,dz=\exp\!\left\{\bm{x}^{\top}\bm{\beta}^{\text{True}}\right\}<+\infty,$
+where $\bm{\beta}^{\text{True}}\in\mathbb{R}^{q+1}$ denotes the true parameter vector.To clarify our strategy, we also recall the definition of estimator consistency, following Casella and Berger (2024).
+An estimator $\widehat{\bm{\beta}}$ of $\bm{\beta}^{\text{True}}$ is said to be consistent (or convergent in probability) if
+$\widehat{\bm{\beta}}\xrightarrow[]{\mathbb{P}}\bm{\beta}^{\text{True}}\quad\text{as }n\to+\infty,$
+that is, if the following condition holds:
+$\lim_{n\to\infty}\Pr\!\left(\|\widehat{\bm{\beta}}-\bm{\beta}^{\text{True}}\|>\epsilon\right)=0,\quad\text{for all }\epsilon>0.$
+4.1 Parameter Vector Estimators Comparison
+Recall that the estimation of the true parameter vector $\bm{\beta}^{\text{True}}$ under the two approaches is carried out by assuming a Tweedie distribution with different weights in each case: the offset approach uses $w=t^{2-p}$, whereas the ratio approach uses $w=t$.
+We denote the corresponding density by $f$, which is defined as
+$f(z\mid\bm{\beta},t)=\exp\!\left(\frac{w}{\phi}\left(\frac{\exp\!\{(1-p)\bm{x}^{\top}\bm{\beta}\}}{1-p}\,z-\frac{\exp\!\{(2-p)\bm{x}^{\top}\bm{\beta}\}}{2-p}\right)+a(z,w,\phi)\right),\qquad z\geq 0,\qquad\bm{\beta}\in\mathbb{R}^{q+1}.$
+We then define the Kullback–Leibler (KL) divergence between the true density $g(\cdot)$ and the density $f(\cdot)$ as
+$KL(\bm{\beta})$ $=$ $\mathbb{E}\!\left[\log\frac{g(z\mid\bm{\beta}^{\text{True}},t)}{f(z\mid\bm{\beta},t)}\right]$
+$=$ $\int_{0}^{+\infty}\log\!\big(g(z\mid\bm{\beta}^{\text{True}},t)\big)\,g(z\mid\bm{\beta}^{\text{True}},t)\,dz$
+$\quad-\int_{0}^{+\infty}\log\!\big(f(z\mid\bm{\beta},t)\big)\,g(z\mid\bm{\beta}^{\text{True}},t)\,dz.$
 Proposition 1.
-The expected value and variance of premium estimator in offset and ratio approach are obtained as follows:
-$\mathrm{E}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]$ $=\exp\left\{\mathbf{X}^{T}_{i}\beta+\frac{\phi}{2}\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}\mathbf{X}_{i}\right\};$ (4.4)
-$\mathrm{E}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right]$ $=\exp\left\{\mathbf{X}^{T}_{i}\beta+\frac{\phi}{2}\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\mathbf{X}_{i}\right\};$ (4.5)
-$\mathrm{Var}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]$ $=\left(\exp\left\{\phi\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}\mathbf{X}_{i}\right\}-1\right)\left(\mathrm{E}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]\right)^{2};$ (4.6)
-$\mathrm{Var}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right]$ $=\left(\exp\left\{\phi\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\mathbf{X}_{i}\right\}-1\right)\left(\mathrm{E}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right]\right)^{2}.$ (4.7)
+According to White (1982), if the Kullback–Leibler divergence between the true density $g$ and the density $f$ (assumed under either the offset or the ratio approach) exists, then the corresponding estimator of the true parameter vector is consistent under both approaches.
+More formally, if
+$KL(\bm{\beta})<+\infty\quad\text{for all }\bm{\beta},$
+then
+$\widehat{\bm{\beta}}^{O}\xrightarrow[]{\mathbb{P}}\bm{\beta}^{\text{True}},\quad\widehat{\bm{\beta}}^{R}\xrightarrow[]{\mathbb{P}}\bm{\beta}^{\text{True}}\quad\text{as }n\to+\infty,$
+where $\bm{\widehat{\beta}}^{O}$ and $\bm{\widehat{\beta}}^{R}$ denote the estimators of $\bm{\beta}^{\text{True}}$ obtained from the offset and ratio approaches, respectively.
 Proof.
-The key to the proof is to note that, for a random variable $Z$, the relationship $\log(Z)\sim\mathcal{N}(\mu,\sigma^{2})$ holds. The expected value and variance of $Z$ are then given by:
-  • $\mathrm{E}\!\left[Z\right]=\exp\{\mu+\frac{1}{2}\sigma^{2}\}$;
-  • $\mathrm{Var}\!\left[Z\right]=\left(\exp\left\{\sigma^{2}\right\}-1\right)\left(\mathrm{E}\!\left[Z\right]\right)^{2}$.
-Additionally, the asymptotic properties of the maximum likelihood estimator (MLE) allow us to derive the distributions of the parameter vector estimators as follows:
-$\bm{\widehat{\beta}^{O}}\sim\mathcal{N}\left(\beta,\phi\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}\right)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \bm{\widehat{\beta}^{R}}\sim\mathcal{N}\left(\beta,\phi\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\right).$
-So, using these asymptotic distributions and Equations (4.2) and (4.2), we obtain the following relationships:
-$\log\left(\bm{\widehat{\zeta}^{O}}_{i}\right)\sim\mathcal{N}\left(\mathbf{X}^{T}_{i}\beta,\phi\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}\mathbf{X}_{i}\right)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \log\left(\bm{\widehat{\zeta}^{R}}_{i}\right)\sim\mathcal{N}\left(\mathbf{X}^{T}_{i}\beta,\phi\mathbf{X}^{T}_{i}\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\mathbf{X}_{i}\right).$
-This demonstrates how the expected values and variances of the premium estimators can be derived using the distributions of the parameter vector estimators. ∎
+The function $KL(\cdot)$ is always nonnegative and finite whenever the probability measures induced by $g(\cdot\mid\bm{\beta}^{\text{True}},t)$ and $f(\cdot\mid\bm{\beta},t)$ are mutually absolutely continuous (see Gray (2011)).
+Since both $g(\cdot\mid\bm{\beta}^{\text{True}},t)$ and $f(\cdot\mid\bm{\beta},t)$ are densities, this condition is equivalent to requiring that their associated probability measures be mutually absolutely continuous, which follows from the Radon–Nikodym theorem.
+Suppose that the minimum of $KL(\bm{\beta})$ exists and denote it by $\bm{\beta}^{*}$.
+According to White (1982), the estimators obtained under the working density $f$ converge in probability to this minimizer $\bm{\beta}^{*}$.
+Therefore, it suffices to show that this minimizer coincides with the true parameter vector, that is,
+$\bm{\beta}^{*}=\bm{\beta}^{\text{True}}.$
+We start from the following equivalence:
+$\arg\min_{\bm{\beta}}KL(\bm{\beta})=\arg\max_{\bm{\beta}}\int_{0}^{+\infty}\log\!\big(f(z\mid\bm{\beta},t)\big)\,g(z\mid\bm{\beta}^{\text{True}},t)\,dz.$
+For notational convenience, we write
+$\int_{0}^{+\infty}\log\!\big(f(z\mid\bm{\beta},t)\big)\,g(z\mid\bm{\beta}^{\text{True}},t)\,dz\;=\;\int_{0}^{+\infty}\log(f)\,g\,dz.$
+Using the expression of the Tweedie density, we obtain
+$\int_{0}^{+\infty}\log(f)\,g\,dz$ $=\int_{0}^{+\infty}\Bigg[\frac{w}{\phi}\Bigg(\frac{\exp\!\{(1-p)\bm{x}^{\top}\bm{\beta}\}}{1-p}\,z-\frac{\exp\!\{(2-p)\bm{x}^{\top}\bm{\beta}\}}{2-p}\Bigg)+a(z,w,\phi)\Bigg]g(z\mid\bm{\beta}^{\text{True}},t)\,dz$
+$\propto\int_{0}^{+\infty}\Bigg(\frac{\exp\!\{(1-p)\bm{x}^{\top}\bm{\beta}\}}{1-p}\,z-\frac{\exp\!\{(2-p)\bm{x}^{\top}\bm{\beta}\}}{2-p}\Bigg)g(z\mid\bm{\beta}^{\text{True}},t)\,dz,$
+where the proportionality follows from the fact that $a(z,w,\phi)$ does not depend on $\bm{\beta}$.
+By linearity of the integral, this expression becomes
+$\frac{\exp\!\{(1-p)\bm{x}^{\top}\bm{\beta}\}}{1-p}\int_{0}^{+\infty}z\,g(z\mid\bm{\beta}^{\text{True}},t)\,dz-\frac{\exp\!\{(2-p)\bm{x}^{\top}\bm{\beta}\}}{2-p}\int_{0}^{+\infty}g(z\mid\bm{\beta}^{\text{True}},t)\,dz.$
+Since
+$\int_{0}^{+\infty}z\,g(z\mid\bm{\beta}^{\text{True}},t)\,dz=\zeta^{\text{True}}=\exp\!\left\{\bm{x}^{\top}\bm{\beta}^{\text{True}}\right\},\qquad\int_{0}^{+\infty}g(z\mid\bm{\beta}^{\text{True}},t)\,dz=1,$
+the objective function can be written as
+$K(\bm{\beta})=\frac{\exp\!\{(1-p)\bm{x}^{\top}\bm{\beta}+\bm{x}^{\top}\bm{\beta}^{\text{True}}\}}{1-p}-\frac{\exp\!\{(2-p)\bm{x}^{\top}\bm{\beta}\}}{2-p}.$
+The gradient of $K(\bm{\beta})$ with respect to $\bm{\beta}$ is
+$\nabla_{\bm{\beta}}K(\bm{\beta})=\Big(\exp\!\{(1-p)\bm{x}^{\top}\bm{\beta}+\bm{x}^{\top}\bm{\beta}^{\text{True}}\}-\exp\!\{(2-p)\bm{x}^{\top}\bm{\beta}\}\Big)\bm{x}.$
+At the minimizer $\bm{\beta}^{*}$, the first-order condition yields
+$\nabla_{\bm{\beta}}K(\bm{\beta}^{*})=\bm{0}\quad\Longleftrightarrow\quad\bm{\beta}^{*}=\bm{\beta}^{\text{True}},$
+which completes the proof. ∎
+4.2 Asymptotic Variance Approximation
+Another important result established by White (1982) is that, under both the offset and ratio approaches, the estimators of the true parameter vector $\bm{\beta}^{\text{True}}$ are asymptotically Gaussian with a common mean equal to $\bm{\beta}^{\text{True}}$.
+Once consistency is established, differences between the two approaches therefore arise solely through their asymptotic dispersion rather than through their limiting value.
+From a practical standpoint, differences in asymptotic dispersion translate into differences in the precision and stability of the estimated covariate effects.
+A smaller asymptotic covariance matrix implies tighter confidence intervals for the regression coefficients, which facilitates the comparison, selection, and validation of pricing variables, particularly when some effects are modest but economically relevant.
+At the same time, reduced dispersion limits the sensitivity of the estimated parameters to random portfolio fluctuations.
+In this context, the comparison of asymptotic covariance matrices provides a natural and meaningful criterion for assessing the relative efficiency of the two estimation strategies.
+Using the asymptotic expansion derived in Appendix Appendix 1: Calculation of Asymptotic Covariance Matrices, we may therefore state the following proposition.
 Proposition 2.
-The expected value of the premium estimator in offset approach is smaller than in the ratio approach.
-Additionally, the variance of the premium estimator in the offset is also smaller than in the ratio approach.
-Formally, for contract $i$, we obtain:
-$\mathrm{E}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]$ $<\mathrm{E}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right];$ (4.8)
-$\mathrm{Var}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]$ $<\mathrm{Var}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right].$ (4.9)
+Consider the leading Gaussian terms of the Edgeworth expansion of the asymptotic covariance matrices of $\widehat{\bm{\beta}}^{O}$ and $\widehat{\bm{\beta}}^{R}$ associated with the offset and ratio approaches.
+Assume that these leading terms are given by
+$\Sigma^{O}$ $=\phi(\bm{X}^{\top}\bm{D}^{O}\bm{X})^{-1},$ (4.1)
+$\Sigma^{R}$ $=\phi(\bm{X}^{\top}\bm{D}^{R}\bm{X})^{-1}.$ (4.2)
+Under the assumptions of Proposition 1 and for risk exposure satisfying $t\leq 1$, these leading Gaussian terms satisfy
+$\Sigma^{R}-\Sigma^{O}\succ 0.$
 Proof.
-Referring to the expressions for the expected values and variances of premium estimators given in Proposition 1, we observe that the key difference between these metrics arises from the asymptotic covariance matrices of the parameter vector estimators in the offset and ratio regressions: $\phi\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1},\phi\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}$.
-This proof, therefore, focuses on comparing these covariance matrices between the two approaches. As noted in Section 2.4.1, both covariance matrices are well defined and positive definite:
-$\phi\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1},\phi\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}\in\mathcal{P}_{q+1}.$
-Let us define the matrix $M$ as the difference between these two covariance matrices:
-$M=\phi\left(\bm{X}^{t}\bm{D}^{R}\bm{X}\right)^{-1}-\phi\left(\bm{X}^{t}\bm{D}^{O}\bm{X}\right)^{-1}.$
-The goal is to show that $M$ is a positive definite matrix.
-To do so, we can utilize the positive definite matrix property described in Section 2.8.
-Therefore, proving that $D^{O}-D^{R}$ is a positive definite matrix is sufficient to show that $M$ is also one.
-By examining Figure 1, we observe that $D^{O}-D^{R}$ is indeed a positive definite matrix. This can be formally expressed as:
-$\forall v=(v_{1},\ldots,v_{n})^{t}\in\mathbf{R}^{n}:v^{t}\left(D^{O}-D^{R}\right)v=\sum_{i=1}^{n}\left(t^{2-p}_{i}-t_{i}\right)\zeta^{2-p}_{i}v^{2}_{i}>0.$
-Additionally, using equation (2.9), we note that the following matrix is positive definite because the matrix $\bm{X}$ has full rank $q+1$:
-$\bm{X}^{t}\left(\bm{D}^{O}-\bm{D}^{R}\right)\bm{X}=\bm{X}^{t}\bm{D}^{O}\bm{X}-\bm{X}^{t}\bm{D}^{R}\bm{X}.$
-Thus $M$ being positive definite implies that the expected value and variance of the premium estimator in the offset approach are smaller than in the ratio approach, as stated in Proposition 2. ∎
-From equations (4.4) and (4.5), it is evident that there is a positive bias in the premium estimator for both regression approaches, as $\mathrm{E}\!\left[\bm{\widehat{\zeta}^{O}}_{i}\right]-\zeta_{i}>0$ and $\mathrm{E}\!\left[\bm{\widehat{\zeta}^{R}}_{i}\right]-\zeta_{i}>0$.
-However, using Proposition 2, we observe that the bias of the premium estimator in the ratio approach is consistently larger than that in the offset approach.
-Therefore, based on the bias and variance criterion, the premium estimator from the offset approach is superior to that of the ratio approach.
-It is important to note that, in insurance pricing, the primary focus is on the bias of the premium estimator.
-Indeed, the goal for an insurer is to develop a pricing model that significantly minimizes premium estimator bias.
-Auto-calibration is generally used for this purpose.
-For further details on premium estimator bias reduction, recent works by Denuit et al.
-(2021), Lindholm et al.
-(2023), Denuit et al.
-(2024) and Denuit and Trufin (2024) can be consulted.
-However, this paper prioritizes bias reduction at the portfolio level rather than at the individual premium level.
-4.3 Random Gaps Comparison
-Another comparison between the two approaches can be made based on what we will refer to as the random gaps.
-More formally, this refers to the following values for the offset and ratio approaches, respectively:
-$\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right).$
-These expressions represent the difference between two random variables: one corresponding to the loss cost of insured individual $i$, and the other to the premium estimator for the same individual.
-We refer to the expectation of the random gap as the "portfolio bias".
-According to Denuit et al.
-(2024), financial equilibrium at the portfolio level corresponds to a zero portfolio bias.
-In other words, the expectation of the random gap should be close to zero to ensure financial equilibrium at the portfolio level.
-However, due to the positive bias in the premium estimators for both regression approaches, financial equilibrium is not achieved in either the offset or ratio approach at the portfolio level.
-Using Proposition 2, we note that the portfolio bias in the ratio approach is smaller than in the offset approach.
-Formally, we establish the following relationship:
-$\mathrm{E}\!\left[\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\right]\leq\mathrm{E}\!\left[\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\right].$ (4.10)
-This relationship implies that the premium estimator from the ratio approach is more effective than that from the offset approach in achieving financial equilibrium across the entire portfolio.
-Our objective is to provide a probabilistic framework for comparing the full distributions of these random gaps using stochastic dominance.
-Specifically, first-order stochastic dominance allows us to demonstrate that the ratio approach results in smaller gaps with higher probability than the offset approach.
-Before presenting our results, we first formally define stochastic dominance between two random variables, based on the work of Shaked and Shanthikumar (2007).
-Definition 4.3.1.
-Consider two random variables, $R_{1}$ and $R_{2}$. We say $R_{2}$ dominates $R_{1}$, denoted by $R_{1}\leq_{st}R_{2}$, if the following condition is satisfied:
-$\mathrm{Pr}\!\left(R_{1}>x\right)\leq\mathrm{Pr}\!\left(R_{2}>x\right),\forall x\in\mathbf{R}.$
-Additionally, we provide a necessary and sufficient condition for ordering two random variables, as outlined in Shaked and Shanthikumar (2007).
-Property 4.3.1.
-For two random variables, R1 and R2, the following relationship holds:
-$R_{1}\leq_{st}R_{2}\Longleftrightarrow\mathrm{E}\!\left[f(R_{1})\right]\leq\mathrm{E}\!\left[f(R_{2})\right],$
-where $f(\cdot)$ is an increasing function for which the expectations exist.
-The two previous results enable us to state the following property:
-Proposition 3.
-The difference between the aggregate loss costs and the sum of the premium estimators is more likely to take a significant value in the offset regression approach than in the ratio regression approach.
-Formally, for contract $i$, we establish the following stochastic dominance:
-$\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\leq_{st}\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right).$ (4.11)
-Proof.
-The key to this proof is using the Taylor series approximation as discussed in Van der Vaart (2000).
-Specifically, the idea is to approximate the expected value of a function of a random variable by applying the function to the expected value of the random variable itself.
-Consider a function $f(\cdot)$.
-By applying the first-order Taylor approximation to the function $f(\cdot)$ at the sums of the difference between the aggregate loss cost and the sum of the premium estimators, $\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)$ and $\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)$, we obtain the following results for each contract $i$:
-$\mathrm{E}\!\left[f\left(\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\right)\right]$ $\approx f\left(\mathrm{E}\!\left[\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\right]\right);$ (4.12)
-$\mathrm{E}\!\left[f\left(\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\right)\right]$ $\approx f\left(\mathrm{E}\!\left[\sum_{i=1}^{n}t_{i}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\right]\right).$ (4.13)
-Using Equation (4.10) and considering $f(.)$ as an increasing function, we conclude that:
-$\mathrm{E}\!\left[f\left(t_{i}\sum_{i=1}^{n}\left(Z_{i}-\bm{\widehat{\zeta}^{R}}_{i}\right)\right)\right]\leq\mathrm{E}\!\left[f\left(t_{i}\sum_{i=1}^{n}\left(Z_{i}-\bm{\widehat{\zeta}^{O}}_{i}\right)\right)\right].$
-This proof shows that the gap between the aggregate loss costs and the sum of the premium estimators is more likely to be larger in the offset regression approach than in the ratio regression approach.
-4.3.1 Remarks
-In general, for each contract $i,i=1,\ldots,n$, when the premium, $\zeta_{i}$, is linked to the score $\mathbf{X}^{T}_{i}\beta$ as $\zeta_{i}=g^{-1}\left(\mathbf{X}^{T}_{i}\beta\right)$, $\widehat{\beta}$ serves as the estimator of $\beta$ in both the offset and ratio approaches.
-However, except in the case of the identity link function, $s>0\mapsto g(s)=s$, the premium estimator $\zeta_{i}$ has non-zero asymptotic bias:
-$\mathrm{E}\!\left[g^{-1}\left(\mathbf{X}^{T}_{i}\widehat{\beta}\right)\right]\neq g^{-1}\left(\mathbf{X}^{T}_{i}\beta\right).$
-Consequently, financial equilibrium based on the expected random gap cannot be achieved unless the identity link is used in premium modelling.
-For this reason, in GLM, insurers consider the balance property to ensure an alternative financial equilibrium based on the alignment of observed loss costs with estimated premiums (see Nelder and Wedderburn (1972) for a discussion on the relationship between the sum of the predicted values and the sum of the observed values in GLM theory).
-As discussed in Section 2, this financial equilibrium ensures that the sum of the estimated premiums equals the aggregate observed loss costs across the entire portfolio, as well as at each level of risk factors used in premium modelling.
-It is important to note that even if the balance property is not verified in either approach, Proposition 3 indicates that it is more likely for the random gaps to take on larger values in the offset approach than in the ratio approach.
-However, this alone is insufficient to determine which approach—offset or ratio—is preferable for achieving financial equilibrium under the balance property.
-Therefore, we need to examine the difference between aggregate observed loss costs and the sum of estimated premiums for both approaches.
-4.4 Observed Gaps Comparison
-We refer to the observed gaps in each regression approach as defined below:
-$\sum_{i=1}^{n}t_{i}\left(z_{i}-\widehat{\zeta}^{R}_{i}\right)\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \sum_{i=1}^{n}t_{i}\left(z_{i}-\widehat{\zeta}^{O}_{i}\right),$ (4.14)
-where $\widehat{\zeta}^{R}_{i}$ and $\widehat{\zeta}^{O}_{i}$ represent the estimated of $\zeta$ from the ratio and offset approaches, respectively.
-These values, $\widehat{\zeta}^{R}_{i}$ and $\widehat{\zeta}^{O}_{i}$ can also be viewed as realizations of the premium estimators in the ratio and offset approaches.
-Our objective is to demonstrate that the observed gaps are equal to zero in the ratio approach when applied to a homogeneous portfolio.
-However, under the same conditions, alignment between aggregate loss costs and the sum of estimated premiums does not occur in the offset approach.
-For a heterogeneous portfolio, the observed gaps in both the offset and ratio approaches are not exactly zero due to the influence of the log-link function in premium modeling.
-Nonetheless, the relationship between the sum of estimated premiums and the aggregate loss costs in the ratio approach can be expressed mathematically.
-This formulation offers valuable insights into the characteristics and magnitude of the observed gap in this approach.
-Our strategy for examining these gaps involves analyzing the MLE for each approach under both homogeneous and heterogeneous portfolio across contracts.
-By maintaining the assumption of $n$ independent contracts and applying consistent notation and assumptions from previous sections, we ensure coherence across models and approaches.
+First, consider $n$ observations of the loss cost per unit of exposure $Z_{1},\ldots,Z_{n}$, with $Z_{i}\sim\operatorname{Tweedie}(\zeta_{i},w_{i},\phi,p)$, $w_{i}\in\{t_{i},t_{i}^{2-p}\}$, and $\zeta_{i}=\exp\{\bm{X}_{i}^{\top}\bm{\beta}\}$.
+According to White (1982), the asymptotic covariance matrix of the estimator is given by $\Sigma=A^{-1}BA^{-1}$, where $A=-I_{n}(\bm{\beta})$ and $B=\mathbb{E}[\nabla(\bm{\beta})\nabla(\bm{\beta})^{\top}]$, with
+$\nabla(\bm{\beta})=\frac{1}{\phi}\Big(\sum_{i=1}^{n}w_{i}\frac{Z_{i}-\zeta_{i}}{\zeta_{i}^{p-1}}x_{i,0},\ldots,\sum_{i=1}^{n}w_{i}\frac{Z_{i}-\zeta_{i}}{\zeta_{i}^{p-1}}x_{i,q}\Big)^{\top}.$
+As shown in Appendix Appendix 1: Calculation of Asymptotic Covariance Matrices, $B=\phi^{-1}\bm{X}^{\top}\bm{D}\bm{X}=-A$, yielding
+$\Sigma=A^{-1}BA^{-1}=-A^{-1}=\phi(\bm{X}^{\top}\bm{D}\bm{X})^{-1}.$
+We now compare the two covariance matrices.
+As noted in Section 2.4.1, $\Sigma^{O},\Sigma^{R}\in\mathcal{P}_{q+1}$.
+Define $M=\Sigma^{R}-\Sigma^{O}$.
+To prove that $M$ is positive definite, it suffices, by the results of Equation (2.8), to show that $\bm{D}^{O}-\bm{D}^{R}$ is positive definite.
+From Figure 1, we have
+$\forall v\in\mathbb{R}^{n},\quad v^{\top}(\bm{D}^{O}-\bm{D}^{R})v=\sum_{i=1}^{n}(t_{i}^{2-p}-t_{i})\zeta_{i}^{2-p}v_{i}^{2}>0.$
+Moreover, since $\bm{X}$ has full rank $q+1$, equation (2.9) implies
+$\bm{X}^{\top}(\bm{D}^{O}-\bm{D}^{R})\bm{X}=\bm{X}^{\top}\bm{D}^{O}\bm{X}-\bm{X}^{\top}\bm{D}^{R}\bm{X}\succ 0,$
+It then follows from Equation (2.8) that $M=\Sigma^{R}-\Sigma^{O}$ is positive definite.
+Taken together, the previous propositions show that each approach can be implemented independently and yields a consistent (i.e., convergent) estimator of the true parameter vector $\bm{\beta}^{\text{True}}$.
+However, based solely on the result stated in Proposition 2, one might be led to conclude that the estimator obtained under the offset approach is asymptotically more efficient than the one obtained under the ratio approach.
+It is important to emphasize that Proposition 2 is formulated exclusively at the level of the leading Gaussian term of the Edgeworth expansion and therefore concerns only first-order asymptotic quantities.
+More specifically, the matrices $\Sigma^{O}$ and $\Sigma^{R}$ represent the leading $O(1/n)$ terms in the covariance expansions of $\widehat{\bm{\beta}}^{O}$ and $\widehat{\bm{\beta}}^{R}$.
+Consequently, the difference $\Sigma^{R}-\Sigma^{O}$ identified in Proposition 2 is itself a first-order quantity of order $O(1/n)$.
+The apparent superiority of the offset approach therefore relies on a comparison between leading terms only and is meaningful insofar as the full asymptotic covariance matrices of $\widehat{\bm{\beta}}^{O}$ and $\widehat{\bm{\beta}}^{R}$ are sufficiently well approximated by their leading Gaussian components $\Sigma^{O}$ and $\Sigma^{R}$, as defined in equations (4.1) and (4.2).
+Since higher-order contributions in the Edgeworth expansion are neglected, the resulting approximation error is of smaller order asymptotically, but may nevertheless be of the same order of magnitude as the first-order difference $\Sigma^{R}-\Sigma^{O}$ in finite samples.
+As a result, it is likely inappropriate to draw definitive conclusions regarding the relative performance of the two approaches based solely on Proposition 2.
+A natural extension would therefore consist in deriving and comparing higher-order terms in the Edgeworth expansion.
+While such an analysis would be of clear theoretical interest, it lies beyond the scope of the present paper.
+4.2.1 Simulations Study
+Instead, we focus on assessing the accuracy of the asymptotic approximation and on comparing the two methods through a simulation study.
+The simulation study is designed solely to assess the empirical relevance of Proposition 2, which provides a first-order asymptotic comparison between the offset and ratio approaches based on the leading Gaussian term of the Edgeworth expansion of the estimators’ covariance matrices.
+To this end, independent samples are generated under a data-generating process that is fully consistent with the modeling assumptions underlying the theoretical result.
+For each scenario, defined by a combination of sample size $n$ and Tweedie power parameter $p$, repeated Monte Carlo replications are performed, and both the offset and ratio estimators are computed using the same Tweedie specification as that used for data generation.
+For each approach, the resulting collection of parameter estimates is used to construct empirical covariance matrices, which provide finite-sample approximations of the true sampling covariances of the estimators.
+The empirical counterpart of the theoretical comparison is then obtained by examining the difference between these covariance matrices and by evaluating its spectral properties, in particular through the sign of its smallest eigenvalue.
+Across all scenarios considered, the empirical results are consistent with the ordering predicted by Proposition 2.
+In particular, the empirical difference $\widehat{\Sigma}^{R}-\widehat{\Sigma}^{O}$ is found to be positive definite in all tested configurations.
+From the standpoint of parameter estimation accuracy, the offset approach may be preferable to the ratio approach in finite samples, as it consistently yields smaller empirical variances for the estimator of $\beta$.
+This empirical finding is in line with the asymptotic ordering established in Proposition 2.
+4.3 Sum of Individual Observed Gaps Comparison
+In the context of automobile insurance pricing, one of the primary objectives is not only parameter estimation but also the accurate determination of premiums.
+From this perspective, we instead focus on proximity to financial balance as a key criterion for evaluating the different approaches.
+As discussed in Denuit et al.
+(2024), financial balance broadly refers to the situation in which the sum of observed total losses coincides with the sum of premiums computed under a given pricing method.
+Since this balance is not exactly achieved under either approach due to the use of a logarithmic link function in premium modeling, financial balance naturally emerges as an important criterion to consider when selecting between the two approaches.
+We consider $n$ independent contracts with observed loss costs $y_{i}=t_{i}z_{i}$ for $i=1,\ldots,n$, where $z_{i}$ denotes the loss cost per unit of risk exposure $t_{i}$.
+We define the individual observed gap as the difference between the observed loss cost and the estimated premium.
+For contract $i$, these gaps are denoted by $\Delta_{i}^{O}$ and $\Delta_{i}^{R}$ under the offset and ratio approaches, respectively:
+$\Delta_{i}^{O}=t_{i}\bigl(z_{i}-\widehat{\zeta}_{i}^{O}\bigr),\qquad\Delta_{i}^{R}=t_{i}\bigl(z_{i}-\widehat{\zeta}_{i}^{R}\bigr),$
+where $\widehat{\zeta}_{i}^{O}$ and $\widehat{\zeta}_{i}^{R}$ are the estimated premiums obtained from the offset and ratio approaches.
+These quantities can be interpreted as realizations of the corresponding premium estimators.
+Our quantities of interest are therefore the aggregated gaps
+$\sum_{i=1}^{n}\Delta_{i}^{O}\quad\text{and}\quad\sum_{i=1}^{n}\Delta_{i}^{R}.$ (4.3)
+Although these sums are generally not equal to zero—reflecting the use of a logarithmic link function rather than the canonical link of the Tweedie model in premium modeling—we show that, under the ratio approach, the sum of individual observed gaps tends to be closer to zero than under the offset approach.
+To analyze this criterion, we rely on the same simulation framework as in the previous section and examine, for each replication with $p=1.5$, the ratio between the total predicted premiums and the total observed losses under both approaches.
+This comparison is illustrated in Figure 2.
+Across all considered sample sizes, the distribution associated with the ratio approach is markedly more concentrated around the value 1—corresponding to a situation close to financial balance—than the distribution obtained under the offset approach.
+In contrast, the offset approach exhibits substantially greater dispersion, reflecting higher variability around the global balance condition, even as the sample size increases.
+A complementary perspective consists in identifying, for each simulation, which approach yields a total predicted premium closer to the observed total losses.
+According to this criterion, and for all scenarios under consideration, the ratio approach is closer to financial balance in approximately 87% of the replications.
+This result further supports the view that, from the standpoint of aggregate financial balance, the ratio approach displays a more stable and robust behavior than the offset approach in finite samples.
+Figure 2: Financial balance density for n=1,000n=1{,}000, 5,0005{,}000, 20,00020{,}000, and 50,00050{,}000 under the offset and ratio approaches.
+Nevertheless, rather than relying solely on simulation results, we argue that additional insight into this financial balance criterion can be gained by examining the parameter estimation equations themselves.
+To this end, we first consider a homogeneous setting in which the sum of individual observed gaps is exactly zero under the ratio approach.
+We then extend the analysis to a heterogeneous setting in order to compare the behavior of this aggregate gap under the ratio and offset approaches.
 4.3.1 Homogeneous Portfolio
@@
 $\ell^{{}^{\prime}}(\zeta)=\frac{1}{\phi\zeta^{p}}\left(\sum_{i=1}^{n}w_{i}z_{i}-\zeta\sum_{i=1}^{n}w_{i}\right).$ (4.5)
-Thus, the MLE of $\zeta$ , denoted by $\widehat{\zeta}$, is obtained as follows: $\widehat{\zeta}=\frac{\sum_{i=1}^{n}w_{i}z_{i}}{\sum_{i=1}^{n}w_{i}}$.
+Thus, the quasi-maximum likelihood estimator (QMLE) of $\zeta$, denoted by $\widehat{\zeta}$, is obtained as follows: $\widehat{\zeta}=\frac{\sum_{i=1}^{n}w_{i}z_{i}}{\sum_{i=1}^{n}w_{i}}$.
   1. 1.
-    Offset Approach: In the offset approach, with $Z_{i}$ as the response variable, the weight for contract $i$ is defined as $w_{i}=t_{i}^{2-p}$.
-Consequently, the MLE for $\zeta$, denoted by $\widehat{\zeta}^{O}$, is:
+    Offset Approach: Under the offset approach, the QMLE of the homogeneous risk parameter $\zeta$ is given by
     $\widehat{\zeta}^{O}=\frac{\sum_{i=1}^{n}t_{i}^{2-p}z_{i}}{\sum_{i=1}^{n}t_{i}^{2-p}}.$ (4.6)
-    The sum of estimated premiums is then calculated as:
-    $\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{O}$ $=\sum_{i=1}^{n}t_{i}\left(\frac{\sum_{i=1}^{n}t_{i}^{2-p}z_{i}}{\sum_{i=1}^{n}t_{i}^{2-p}}\right)$
-    $=\frac{\sum_{i=1}^{n}t_{i}^{2-p}z_{i}}{\sum_{i=1}^{n}t_{i}^{2-p}}\sum_{i=1}^{n}t_{i}.$
-    We observe that the estimated premiums sum matches the aggregate observed loss cost only under the conditions that $t_{i}=1$ or $p=1$.
-Given our assumption that the variance parameter $21<p<$ 2 and $t_{i}<1$, it follows that:
-    $\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{O}\neq\sum_{i=1}^{n}t_{i}z_{i}.$
+    The corresponding aggregate estimated premium is therefore
+    $\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{O}=\left(\sum_{i=1}^{n}t_{i}\right)\frac{\sum_{i=1}^{n}t_{i}^{2-p}z_{i}}{\sum_{i=1}^{n}t_{i}^{2-p}}.$ (4.7)
+    Exact financial balance, i.e.
+    $\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{O}=\sum_{i=1}^{n}t_{i}z_{i},$
+    holds if and only if the $t_{i}^{2-p}$-weighted average of the observed loss costs $z_{i}$ coincides with their $t_{i}$-weighted average.
+This condition is automatically satisfied in degenerate cases such as $p=1$, or when all exposures $t_{i}$ are equal, in which case the two weighting schemes coincide.
+However, under the general setting considered in this paper, where $1<p<2$ and exposures satisfy $t_{i}\leq 1$ with heterogeneous durations, there is no structural mechanism in the offset model that enforces this identity.
+As a result, aggregate financial balance is not guaranteed under the offset approach and may fail in general, although it can still occur for particular realizations of the portfolio.
+Consequently, the aggregate deviation
+    $\sum_{i=1}^{n}\Delta_{i}^{O}=\sum_{i=1}^{n}t_{i}\widehat{\zeta}^{O}-\sum_{i=1}^{n}t_{i}z_{i}$
+    is not constrained to vanish.
   2. 2.
     Ratio Approach: In the ratio approach, where $Z_{i}$ also serves as the response variable, the weight for contract $i$ is defined as $w_{i}=t_{i}$.
-The MLE for $\zeta$, denoted by $\widehat{\zeta}^{R}$, is calculated as:
+The QMLE for $\zeta$, denoted by $\widehat{\zeta}^{R}$, is calculated as:
     $\widehat{\zeta}^{R}=\frac{\sum_{i=1}^{n}t_{i}z_{i}}{\sum_{i=1}^{n}t_{i}}.$ (4.8)
@@
     $=\sum_{i=1}^{n}t_{i}z_{i}.$
+    So, we conclude that $\sum_{i=1}^{n}\Delta_{i}^{R}=0$.
 This result suggests that the ratio approach inherently ensures that the sum of estimated premiums aligns with the aggregate observed loss costs, making it advantageous for achieving balance at the portfolio level.
 4.3.2 Heterogeneous Portfolio
-We now compare the approaches in the case of a heterogeneous portfolio with explanatory variables.
-We then incorporate the characteristics vector $\mathbf{X}_{i}$ to model the premium for each contract $i$, allowing for contract-specific premium estimation and accounting for variation across contracts.
-Thus, the premium in this approach is represented by Equation (4.1).
-Unlike the homogeneous portfolio, where a closed-form solution for the MLE of $\beta$ was feasible, here the MLE for the parameter vector $\beta$ does not have a closed-form solution due to its dependence on $\mathbf{X}_{i}$.
-Consequently, we apply the IRLS algorithm (see Section 2.4.1) to obtain the MLE of $\beta$ for both the offset and ratio approaches, as discussed in previous sections.
-To better understand the gap between the aggregate observed loss costs and the sum of estimated premiums in each approach, we initialize the IRLS algorithm with values of $\beta$ based on the homogeneous portfolio assumption.
+The homogeneous framework can be extended to account for heterogeneity by incorporating the covariate vector $\mathbf{X}_{i}$ in the premium model for each contract $i$.
+This allows for contract-specific premium estimation and captures variation across contracts.
+In contrast to the homogeneous portfolio, where a closed-form expression for the QMLE of $\bm{\beta}$ was available, the presence of covariates implies that the QMLE of the parameter vector $\bm{\beta}$ no longer admits a closed-form solution.
+This is due to its dependence on $\mathbf{X}_{i}$.
+Consequently, the QMLE of $\bm{\beta}$ is obtained using the IRLS algorithm (see Section 2.4.1) for both the offset and ratio approaches, as discussed in the previous sections.
+To better understand the gap between the aggregate observed loss costs and the sum of estimated premiums in each approach, we initialize the IRLS algorithm with values of $\bm{\beta}$ based on the homogeneous portfolio assumption.
 This initialization promotes consistency in comparing the offset and ratio approaches and clarifies how incorporating $\mathbf{X}_{i}$ affects the alignment of estimated premiums with aggregate loss costs.
@@
     Offset Approach: We initialize the IRLS algorithm with:
-    $\beta_{0}=\log\left(\widehat{\zeta}^{O}\right),\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \beta_{j}=0,\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ j=1,\ldots,q.$
-    By denoting the initial value by $\beta^{O}_{(0)}$, the estimate of the parameter vector at iteration $K$, $\beta^{O}_{(K)}$, is obtained as follows:
-    $\beta^{O}_{(K)}=K\beta^{O}_{(0)}+\underbrace{\sum_{k=1}^{K-1}\left(I^{O}_{n}\left(\beta^{O}_{(k)}\right)\right)^{-1}\nabla^{O}\left(\beta^{O}_{(k)}\right)}_{\delta^{O}_{K-1}}=K\beta^{O}_{(0)}+\delta^{O}_{K-1},$
+    β_0 = log(^ζ^O), and β_j = 0, for j=1,…,q.
+    By denoting the initial value by $\bm{\beta}^{O}_{(0)}$, the estimate of the parameter vector at iteration $K$, $\bm{\beta}^{O}_{(K)}$, is obtained as follows:
+    β^O_(K) = K β^O_(0) + ⏟∑_k=1^K - 1 (I^O_n(β^O_(k)))^-1 ∇^O(β^O_(k) )_δ^O_K - 1 = K β^O_(0) + δ^O_K - 1,
     where $I^{O}_{n}(\cdot)$ is the Fisher information matrix for the offset approach.
 Therefore, we calculate the estimate premium at iteration $K$, denoted by $\widehat{\zeta}^{O}_{i(K)}$, as follows:
-    $\widehat{\zeta}^{O}_{i(K)}=\left(\widehat{\zeta}^{O}\right)^{K}\exp\left\{\mathbf{X}^{T}_{i}\delta^{O}_{K-1}\right\},$
+    ^ζ^O_i(K) = (^ζ^O)^K exp{ X^T_i δ^O_K - 1},
     where $\widehat{\zeta}^{O}$, defined by equation (4.6), represents the annual premium for all insureds in the offset approach for the homogeneous portfolio.
@@
     Ratio Approach: We initialize the IRLS algorithm with:
-    $\beta_{0}=\log\left(\widehat{\zeta}^{R}\right),\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \beta_{j}=0,\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ j=1,\ldots,q.$
-    By denoting the initial value by $\beta^{R}_{(0)}$, the estimate of the parameter vector at iteration $K$, $\beta^{R}_{(K)}$, is obtained as follows:
-    $\beta^{R}_{(K)}=K\beta^{R}_{(0)}+\underbrace{\sum_{k=1}^{K-1}\left(I^{R}_{n}\left(\beta^{R}_{(k)}\right)\right)^{-1}\nabla^{R}\left(\beta^{R}_{(k)}\right)}_{\delta^{R}_{K-1}}=K\beta^{R}_{(0)}+\delta^{R}_{K-1},$
+    β_0 = log(^ζ^R), and β_j = 0, for j=1,…,q.
+    By denoting the initial value by $\bm{\beta}^{R}_{(0)}$, the estimate of the parameter vector at iteration $K$, $\bm{\beta}^{R}_{(K)}$, is obtained as follows:
+    β^R_(K) = K β^R_(0) + ⏟∑_k=1^K - 1 (I^R_n(β^R_(k)))^-1 ∇^R(β^R_(k) )_δ^R_K - 1 = K β^R_(0) + δ^R_K - 1,
     where $I^{R}_{n}(\cdot)$ is the Fisher information matrix for the ratio approach. Therefore, we calculate the estimate premium at iteration $K$ as follows:
-    $\widehat{\zeta}^{R}_{i(K)}=\left(\widehat{\zeta}^{R}\right)^{K}\exp\left\{\mathbf{X}^{T}_{i}\delta^{R}_{K-1}\right\},$
+    ^ζ^R_i(K) = ( ^ζ^R)^K exp{ X^T_i δ^R_K - 1},
     where $\widehat{\zeta}^{R}$, defined by equation (4.8), represents the annual premium for all insureds in the ratio approach for the homogeneous portfolio.
@@
 4.3.3 Numerical Illustration
-To further compare the observed gaps in both regression approaches, especially for the heterogeneous portfolio where an explicit analytical solution is not available, we conduct a numerical example.
-The following assumptions are made:
+To better understand the implications of the parameter estimation equations underlying the two approaches—particularly in the heterogeneous case, where no explicit analytical solution is available—we propose a simple numerical illustration.
+The objective of this illustration is not to assess the statistical performance of the estimators, but rather to highlight the structural differences between the estimation equations and to clarify how these differences translate into distinct behaviors with respect to the financial balance criterion.
+The setup is as follows:
   1. 1.
-    Portfolio Setup: In the homogeneous case, we consider $n=100$ independent contracts with no covariates.
-In the heterogeneous case, we include two risk factors in the mean parameter of each contract, generated from binomial distributions:
-    • $\mathbf{x_{1}}\sim\text{Binomial}(100,0.75)$
-    • $\mathbf{x_{2}}\sim\text{Binomial}(100,0.15)$
+    Portfolio setup.
+In the homogeneous case, we consider a portfolio of $n=100$ independent contracts without covariates.
+In the heterogeneous case, two risk factors are incorporated into the mean specification of each contract.
+These covariates are generated independently from binomial distributions:
+    • $x_{1}\sim\text{Binomial}(100,0.75)$,
+    • $x_{2}\sim\text{Binomial}(100,0.15)$.
+    This specification induces heterogeneity in the portfolio while remaining sufficiently simple to allow a transparent interpretation of the estimation equations.
   2. 2.
-    Risk Exposure: Risk exposures are randomly drawn from a uniform distribution over the interval $\left[\frac{30}{365},\frac{335}{365}\right]$, and then sorted in increasing order to assign each contract a corresponding rank $i$.
+    Risk exposure. Exposure levels $t_{i}$ are drawn from a uniform distribution on the interval
+    $\left[\tfrac{30}{365},\tfrac{335}{365}\right],$
+    and then sorted in increasing order to assign contract indices $i=1,\ldots,n$.
+This ordering reflects a common actuarial situation in which higher indexed contracts correspond to larger exposure levels.
   3. 3.
-    Loss Cost Scenarios:
-    • Increasing Scenario: Loss costs increase with risk exposure: $y_{i}=i$, for $i=1,\ldots,n$.
-    • Decreasing Scenario: Loss costs decrease with risk exposure: $y_{i}=n-i-1$, for $i=1,\ldots,n$.
-This small numerical example allows us to highlight some of the analytical results discussed earlier.
-More importantly, it helps to better understand the behavior of the observed gaps for a heterogeneous portfolio, which may not always be expressed analytically based on the approaches (as discussed in Section 4.4.2).
-Thus, for the Increasing and the Decreasing scenarios, we define the individual gap for contract $i$ using the following equations:
-$\Delta_{i}^{R}=t_{i}\left(z_{i}-\widehat{\zeta}^{R}_{i}\right)\quad\text{and}\quad\Delta_{i}^{O}=t_{i}\left(z_{i}-\widehat{\zeta}^{O}_{i}\right).$
-Both $\Delta_{i}^{R}$ and $\Delta_{i}^{O}$ were already used in equation (4.14).
-For risk exposure in $\left[\frac{30}{365},\frac{335}{365}\right]$, the differences $\Delta_{i}$ are plotted as shown in Figure 2.
-Depending on the portfolio (homogeneous or heterogeneous), the approach (offset or ratio), and the scenario (increasing or decreasing) being studied, many conclusions can be made from Figure 2.
-  • Ratio Approach (with solid and dashed blue lines)
-    We begin the analysis by recalling that, as shown in Section 4.4.1, $\sum_{i=1}^{n}\Delta_{i}^{R}=0$ for the homogeneous portfolio.
-This corresponds to the blue dashed line in both graphs of Figure 2.
-In comparison, the two curves (solid blue lines) representing the heterogeneous portfolio case are very close to the homogeneous case, indicating that $\sum_{i=1}^{n}\Delta_{i}^{R}$ for the heterogeneous case is expected to be near zero.
-However, the small jumps observed in both solid blue lines suggest that full financial equilibrium cannot be fully achieved.
-  • Offset Approach (with solid and dashed red lines)
-    We showed in Section 4.4.1 that the offset approach does not allow for $\sum_{i=1}^{n}\Delta_{i}^{O}=0$, which is exactly what we observe with all the red lines of the homogeneous portfolio.
-Since the sum of the values of the dashed blue line equals zero, the upward (or downward) shift of the red line indicates that financial equilibrium is not achieved.
-We also observe the same small jumps for the solid blue line associated with the heterogeneous portfolio.
-In the scenario where costs increase with exposure, it is interesting to note that the offset approach line is shifted upwards compared to the ratio approach line, while the opposite shift is observed in the decreasing scenario.
-Thus, we can see that the gap in financial equilibrium for the offset approach will be either positive or negative, depending on whether we are in an increasing or decreasing cost-exposure scenario.
-Figure 2: Observed gap by risk exposure (left: Scenario 1, right: Scenario 2)
+    Loss cost. Rather than assuming a stochastic loss-generating mechanism, we consider a deterministic specification given by
+    $y_{i}=i,\qquad i=1,\ldots,n.$
+    This choice is deliberate and aims to isolate the effect of the estimation equations themselves.
+By eliminating random noise, the resulting comparison focuses exclusively on the structural differences between the offset and ratio approaches and on their implications for aggregate financial balance.
+Figure 3: Individual observed gaps: single heterogeneous realization (left) and across multiple heterogeneous realizations (right).
+We then apply the corresponding parameter estimation equations to this portfolio.
+For a single realization, the individual observed gaps are displayed in the left panel of Figure 3.
+This panel allows a direct comparison across portfolio types (homogeneous versus heterogeneous) and estimation approaches (offset versus ratio).
+Several qualitative insights can be drawn from this figure:
+  • Ratio approach (solid and dashed blue curves).
+    Recall that, for a homogeneous portfolio, the ratio approach satisfies the exact aggregate balance condition $\sum_{i=1}^{n}\Delta_{i}^{R}=0$, as established in Section 4.3.1.
+Although this aggregate quantity is not explicitly reported in the figure, the dashed blue curve corresponds to the individual observed gaps obtained under the homogeneous assumption and therefore sums to zero by construction.
+    This homogeneous benchmark provides a natural reference for interpreting the solid blue curve, which represents the individual observed gaps obtained under the heterogeneous specification.
+The two curves are remarkably close over the entire range of exposure levels, indicating that the introduction of covariates does not materially alter the aggregate balance property of the ratio approach.
+In particular, this visual proximity suggests that the sum $\sum_{i=1}^{n}\Delta_{i}^{R}$ in the heterogeneous case remains close to zero.
+    At the same time, the small local deviations and step-like variations observed along the solid blue curve reflect the fact that exact financial balance cannot generally be achieved in the heterogeneous setting, except in accidental cases where the sum of these local variations happens to cancel out.
+These discrepancies arise from the contract-specific nature of the estimated premiums and from the iterative nature of the estimation procedure.
+Nevertheless, the overall behavior remains strongly anchored to the homogeneous equilibrium, highlighting the robustness of the ratio approach with respect to portfolio heterogeneity.
+  • Offset approach (solid and dashed red curves).
+    As established in Section 4.3.1, the offset approach does not satisfy an exact aggregate balance condition, that is, $\sum_{i=1}^{n}\Delta_{i}^{O}\neq 0.$ This feature is directly visible in Figure 3, where both red curves display a systematic displacement relative to the ratio approach, which serves as a natural benchmark since its aggregate gap is zero by construction.
+    In the homogeneous case (dashed red curve), this displacement is already apparent, as the offset curve lies uniformly above its ratio counterpart, indicating a departure from financial equilibrium.
+In the heterogeneous case (solid red curve), the offset curve remains centered around its homogeneous counterpart, suggesting that the imbalance persists even in the presence of additional local variability arising from contract-specific covariates.
+To further assess the impact of parameter estimation on financial balance, we consider a Monte Carlo experiment based on 1,000 heterogeneous portfolios generated from varying realizations of $x_{1}$ and $x_{2}$.
+The corresponding results are reported in the right panel of Figure 3.
+In this panel, all 1,000 individual gap curves are displayed simultaneously as semi-transparent lines, providing a direct visualization of the outcomes associated with heterogeneous portfolios.
+To summarize this variability, a pointwise 95% envelope is added in the form of a shaded ribbon, while the bold solid curve represents the median gap across replications.
+We can see that the individual gap curves obtained across replications are strongly concentrated around their respective homogeneous benchmarks for both approaches.
+This is reflected by the close proximity of the median heterogeneous curve to the corresponding homogeneous curve shown in the left panel, as well as by the relatively narrow width of the 95% ribbons over the entire range of risk exposure.
+These patterns indicate that heterogeneity primarily introduces dispersion around the homogeneous structure, rather than inducing systematic shifts in the gap curves themselves.
+An important implication of this concentration phenomenon is that the aggregate imbalance observed under the homogeneous specification provides a meaningful first-order diagnostic for the heterogeneous case.
+Since heterogeneity mainly adds random fluctuations around the homogeneous benchmark, a substantial lack of financial balance at the homogeneous level is unlikely to be systematically corrected once heterogeneity is introduced.
+Conversely, when the homogeneous imbalance is small, near-balance under heterogeneous modeling becomes more plausible, albeit still not structurally enforced.
+In this respect, the two approaches differ fundamentally.
+Exact aggregate balance holds by construction for the homogeneous ratio approach, whereas no such property holds for the homogeneous offset approach.
+As a result, the concentration of heterogeneous gap curves around their homogeneous counterparts implies that, in heterogeneous portfolios, the ratio approach is typically closer to financial equilibrium than the offset approach.
+Finally, we note that the offset approach may occasionally exhibit a better financial balance than the ratio approach for specific realizations, as observed in the simulation results reported in Figure 2, where it is selected in approximately 13% of the replications.
+As illustrated in Figure 3, such outcomes can partly be attributed to more pronounced random fluctuations, whose aggregate effect may incidentally compensate for the residual imbalance.
+However, these situations are also more likely to occur in scenarios where the homogeneous offset specification itself happens to be close to financial balance.
+In such cases, the additional dispersion introduced by heterogeneity may suffice to produce a near-balanced aggregate outcome under the offset approach.
+Since this compensation mechanism is not structural but contingent on the particular data realization and on the initial homogeneous imbalance, the ratio approach generally provides a more reliable and robust financial balance across heterogeneous portfolios.
 5 Empirical Illustration
@@
 As such, we believe the resulting pricing model reflects standard industry practices.
-Descriptive statistics for these covariates are provided in the appendix.
+Descriptive statistics for these covariates are provided in Figure 6.
+Figure 6: Descriptive statistics of all 8 covariates from the database
 5.2 Offset and Ratio Approaches
@@
 5.2.1 Estimated Parameters
-The estimated parameters $\beta$ for covariates $X_{1}-X_{8}$ can be different for both approaches.
+The estimated parameters $\bm{\beta}$ for covariates $X_{1}-X_{8}$ can be different for both approaches.
 To compare the estimated values, we compute a coefficient ratio for each covariate by dividing the estimated coefficient from the offset approach by the corresponding coefficient from the ratio approach.
@@
 These results are displayed in Figure 7.
-Without suprise, the estimated coefficients are identical in both the offset and ratio approaches, which was expected because the offset and ratio approaches yield equivalent results when the risk exposure is equal to 1.
+Without surprise, the estimated coefficients are identical in both the offset and ratio approaches, which was expected because the offset and ratio approaches yield equivalent results when the risk exposure is equal to 1.
 On the other hand, for contracts with mid-term cancellations, the estimated coefficients differ noticeably between the offset and ratio approaches.
 This divergence is anticipated because, as discussed in Section 3, the log-likelihood and gradient functions differ for the two approaches when risk exposure is less than 1.
-The significant differences, particularly in coefficients associated with covariates $X1$, $X2$, and $X4$, highlight how the choice of approach impacts premium estimation.
+The significant differences, particularly in coefficients associated with covariates $X_{1}$, $X_{2}$, and $X_{4}$, highlight how the choice of approach impacts premium estimation.
 These differences imply that premiums for certain policyholder profiles will vary substantially based on the chosen approach.
@@
 These ratios are plotted against aggregate observed loss costs in Figure 9, with risk classes labeled as integers from 1 to 17.
-In the Full exposure group, the observed gap is identical for both the offset and ratio approaches, explained by the identical estimated premiums in these methods.
-As previously noted, particularly in Section 4.3.1, it is interesting to observe that the sums of predictions and observations by risk class differ even when risk exposure equals one.
-In the Mid-term cancellation group, however, some differences emerge in the observed gaps between approaches.
-Here, the ratio approach appears to perform better, with the sum of estimated premiums aligning more closely with the aggregate observed loss costs than in the offset approach.
-Figure 7: Observed gaps comparison (left: Full exposure, right: Mid-term cancellation)
+In the full-exposure group, the sum of individual observed gaps is identical for both the offset and ratio approaches, which is explained by the fact that the estimated premiums are identical under the two methods.
+As noted previously, it is nevertheless interesting to observe that, in the heterogeneous data, the sums of predicted and observed losses by risk class may differ even when the risk exposure equals one.
+A similar phenomenon is observed under homogeneous data when the offset approach is used.
+For the mid-term cancellation group, all ratios are reported in the right panel of Figure 9.
+The ratio approach yields a closer alignment between the sum of estimated premiums and the aggregate observed loss costs than the offset approach.
+In particular, the ratio of the sum of estimated premiums to the aggregate observed loss cost under homogeneous data using the offset approach is equal to $1.22$, whereas the ratio approach yields exact aggregate balance under homogeneity.
+As discussed in Section 4.3.3, ratios of similar magnitude are observed for both approaches in the heterogeneous case.
+Figure 9: Sum of individual observed gaps comparison (left: Full exposure, right: Mid-term cancellation)
 5.3 Discussion
 Even though neither of the two approaches studied is perfect in every respect, the advantages of using the ratio approach for a Tweedie regression are clear for the insurance portfolio we analyzed.
-However, we also recognize that the analysis of the approach to be taken has brought to light several important issues, which may lead to different choices or potential solutions depending on the context:
-  1. 1.
-    Approach Choice and Portfolio Composition:
-    As shown in Figure 1, the influence of the chosen approach is relatively minor for portfolios with few mid-term cancellations, such as home insurance or certain commercial insurance lines.
-In these cases, where most contracts have full-year exposures, the offset and ratio approaches produce similar results.
-However, for insurance portfolios with a higher proportion of contracts having partial-year exposures — typical in certain types of insurance, such as North American car insurance — the choice between the offset and ratio approaches becomes critical.
-  2. 2.
-    Impact on Premiums and Policyholder Fairness:
-    The different coefficient estimates produced by the offset and ratio approaches imply that premiums for the same insured can vary significantly depending on the approach (see the value of estimated parameters from Figure 5).
-This difference of premium raises important considerations in a competitive insurance market, as it could influence which policyholder profiles benefit or are disadvantaged by each approach.
-Understanding how these differences impact policyholder fairness and market competitiveness could help insurers decide which premium structure better aligns with market expectations and policyholder satisfaction.
-  3. 3.
-    Mid-term Cancellations and Loss Costs:
-    Table 1 and Figure 3 demonstrate that mid-term cancellations significantly affect loss costs, highlighting the importance of addressing cancellations more precisely in ratemaking models.
-Rather than simply incorporating risk exposure within a single pricing model, as done with a Tweedie distribution in this paper, analyzing specific causes of cancellations could enable the development of more tailored approaches.
-By examining the reasons behind cancellations, insurers could create models that account for different cancellation behaviours and adjust premiums accordingly, potentially leading to more accurate pricing strategies.
+However, the analysis of the modeling choices considered in this paper also brings to light several important issues, which may lead to different conclusions or modeling strategies depending on the portfolio structure and the underlying risk characteristics.
+First, the relevance of the offset or ratio approach strongly depends on the composition of the insurance portfolio.
+As illustrated in Figure 1, the impact of the chosen approach remains relatively limited for portfolios characterized by a low frequency of mid-term cancellations, such as home insurance or certain commercial insurance lines, where most contracts exhibit full-year exposures.
+In such contexts, the offset and ratio approaches yield similar results.
+In contrast, for portfolios with a substantial proportion of partial-year exposures—typical of certain insurance markets, such as North American automobile insurance—the choice between the two approaches becomes critical and can materially affect the modeling outcomes.
+These differences are not only technical but also have direct implications for premium determination and policyholder fairness.
+As shown by the estimated parameters reported in Figure 7, the offset and ratio approaches may lead to substantially different premiums for the same insured risk.
+In a competitive insurance market, such discrepancies may influence which policyholder profiles are favored or disadvantaged by each approach.
+Understanding how these pricing differences affect fairness perceptions and market competitiveness is therefore an important consideration when selecting a ratemaking methodology.
+The empirical results further emphasize the central role of mid-term cancellations in shaping loss costs.
+Table 1 and Figure 4 show that cancellations have a significant impact on observed loss costs, underlining the need to address this phenomenon more explicitly in pricing models.
+Rather than incorporating exposure mechanically within a single Tweedie framework, as done in this paper, a more detailed analysis of cancellation behavior could allow insurers to develop more refined modeling strategies.
+In particular, distinguishing between different causes of cancellations may help capture heterogeneous risk dynamics and improve premium adequacy.
+Finally, the analysis raises broader questions regarding the assumption of strict linear proportionality between exposure and expected loss cost, an assumption that underlies both the offset and ratio approaches.
+While this assumption is widely adopted in actuarial practice, the descriptive results presented in Figure 5 suggest that loss costs do not necessarily scale linearly with exposure.
+Shorter exposure periods, in particular, appear to be associated with systematically different loss dynamics, consistent with several scenarios discussed in Section 5.1.1.
+Although both approaches considered here impose linearity by construction, these empirical patterns indicate that relaxing this assumption could provide additional modeling flexibility and lead to a more accurate representation of exposure-related heterogeneity.
 6 Conclusion
-This study compares the ratio and offset approaches to premium estimation, focusing on how each method incorporates risk exposure.
-The ratio approach assumes a contract’s premium is directly proportional to its risk exposure, treating the exposure as a weight independent of the response variable’s distribution.
-In contrast, the offset approach, particularly with a logarithmic link function, incorporates the log of the risk exposure as an offset variable in the premium model.
-Our findings show that these approaches are not equivalent when the response variable follows a Tweedie distribution.
-Through a detailed comparison, we identified the respective strengths and weaknesses of each approach.
-Additionally, we demonstrated that the offset approach can be reformulated as a weighted regression, structurally similar to the ratio approach.
-Thus, the distinction between these methods primarily lies in the choice of weights, framing the comparison as one between two weighted regressions.
-Our analysis of weight selection revealed that, for any given contract, the weight in the offset approach is consistently greater than that in the ratio approach.
-This supports our conclusion that the premium estimator from the offset approach generally outperforms the ratio approach, especially when considering bias and variance minimization.
-However, our assessment of the random gap between aggregate loss costs and the sum of premium estimators suggests that this gap tends to be larger for the offset approach than for the ratio approach.
-To evaluate the effectiveness of each approach for pricing, we assessed the balance property, which aligns the sum of estimated premiums with the aggregate actual loss costs.
-Based on this criterion, the ratio approach performed better than the offset approach.
-As discussed in Section 5.3, we recommend the ratio approach, particularly in portfolios with diverse risk exposures, due to its ability to better align premiums with actual loss costs.
-For future research, comparing the premium estimators of both methods using auto-calibration techniques, as suggested by Denuit et al.
-(2021), Lindholm et al.
-(2023), Denuit et al.
-(2024) and Denuit and Trufin (2024), would be valuable, especially given the positive bias observed in both approaches.
-Appendix: Descriptive statistics of covariates
-Figure 8: Descriptive statistics of all 8 covariates from the database
+This paper has examined two widely used strategies for incorporating risk exposure into premium estimation under Tweedie regression models: the offset approach and the ratio approach.
+Although both methods rely on the same proportionality assumption between exposure and expected loss, we have shown that they induce fundamentally different weighting structures and therefore lead to distinct statistical and financial properties.
+From a purely inferential perspective, both approaches yield consistent estimators of the true parameter vector under mild regularity conditions.
+A first-order asymptotic analysis based on the leading Gaussian term of the Edgeworth expansion suggests that the offset approach may exhibit greater efficiency, in the sense of a smaller asymptotic covariance matrix.
+Simulation results seem to confirm this comparison, indicating that in finite samples the offset approach also displays smaller empirical variance.
+Beyond parameter estimation, this paper emphasizes that premium modeling in insurance is inherently a financial exercise, for which aggregate balance considerations play a central role.
+From this standpoint, a key contribution of this work is to show that the ratio approach exhibits a structurally stronger proximity to financial equilibrium.
+In homogeneous portfolios, it satisfies an exact aggregate balance condition by construction, while in heterogeneous settings it remains systematically closer to balance than the offset approach.
+Importantly, this property is not driven by asymptotics or incidental compensation effects, but follows directly from the structure of the estimating equations.
+In the final section of the paper, we complement the theoretical and simulation-based analyses with an empirical study based on a real Canadian automobile insurance portfolio.
+This application allows for a direct comparison of the offset and ratio approaches both in terms of parameter estimates—where the two approaches yield different values—and in terms of aggregate financial balance.
+The empirical results indicate that the ratio approach achieves a closer alignment between the sum of estimated premiums and the aggregate observed loss costs, whereas the offset approach exhibits a systematic imbalance.
+The empirical results also suggest that a more refined examination of the relationship between risk exposure and observed loss costs may be warranted.
+In particular, the assumption of a strictly linear proportionality between exposure and expected loss appears to be restrictive in the presence of mid-term cancellations and heterogeneous exposure patterns, even though such proportionality is implicitly imposed by both modeling approaches considered in this paper.
+One important source of this heterogeneity arises when vehicles are replaced during the policy term, either following the sale and purchase of another vehicle or, as discussed in Section 5, after a significant loss that renders the original coverage unnecessary.
+In this context, more detailed tracking of vehicle replacements within insurers’ databases could help improve the representation of exposure-related risk.
+More generally, these findings indicate that, while risk exposure $t$ remains a vehicle-level quantity, its relationship with risk may depend on the broader exposure configuration of the policy to which the vehicle belongs.
+In particular, when multiple vehicles are insured sequentially or concurrently under the same policy, considering interactions among vehicle-level exposures may improve the representation of the underlying risk structure.
+Appendix 1: Calculation of Asymptotic Covariance Matrices
+This appendix summarizes results from White (1982) that are directly relevant to Proposition 2.
+We provide additional details on the calculation of the asymptotic covariance matrices of the parameter estimators in the Offset and Ratio approaches.
+To this end, we begin by recalling the expression of the response variance in both approaches, which is fundamental for deriving the covariance matrices:
+$\mathrm{Var}\!\left[Z_{i}\mid\mathbf{X}_{i},w_{i}\right]=\frac{\phi}{w_{i}}\zeta_{i}^{p},\quad i=1,\ldots,n,$
+where $\zeta_{i}=\exp\{\mathbf{X}_{i}^{\top}\bm{\beta}\}$, and $w_{i}=t_{i}^{2-p}$ under the Offset approach while $w_{i}=t_{i}$ under the ratio approach.
+For convenience, we also introduce the notation
+$\bm{D}(W)=\operatorname{diag}\!\left(w_{i}\zeta_{i}^{2-p}\right)_{i=1,\ldots,n},$
+with the vector $\bm{W}=(w_{1},\ldots,w_{n})^{\top}$.
+This expression for the variance is used in the computation of the asymptotic covariance matrix of the parameter estimator.
+Following White (1982), this matrix is defined as
+$\Sigma=A^{-1}BA^{-1},$
+where $A=-\tfrac{1}{\phi}\bm{X}^{\top}\bm{D}(W)\bm{X}$ and $B=\mathbb{E}\!\left[\nabla(\bm{\beta})\nabla(\bm{\beta})^{\top}\right]$, with the score vector
+$\nabla(\bm{\beta})=\frac{1}{\phi}\sum_{i=1}^{n}w_{i}\frac{Z_{i}-\zeta_{i}}{\zeta_{i}^{p-1}}\mathbf{X}_{i}.$
+Then, the matrix $B$ is
+$B$ $=\frac{1}{\phi^{2}}\mathrm{E}\!\left[\left(\sum_{i=1}^{n}w_{i}\frac{Z_{i}-\zeta_{i}}{\zeta_{i}^{p-1}}x_{i,j_{1}}\right)\left(\sum_{i=1}^{n}w_{i}\frac{Z_{i}-\zeta_{i}}{\zeta_{i}^{p-1}}x_{i,j_{2}}\right)\right]_{j_{1},j_{2}=0,\ldots,q}$
+$=\frac{1}{\phi^{2}}\mathrm{E}\!\left[\sum_{i=1}^{n}w_{i}^{2}\frac{(Z_{i}-\zeta_{i})^{2}}{\zeta_{i}^{2p-2}}x_{i,j_{1}}x_{i,j_{2}}\right]_{j_{1},j_{2}=0,\ldots,q}$
+$=\frac{1}{\phi^{2}}\sum_{i=1}^{n}w_{i}^{2}\frac{\mathrm{Var}\!\left[Z_{i}\mid\mathbf{X}_{i},w_{i}\right]}{\zeta_{i}^{2p-2}}x_{i,j_{1}}x_{i,j_{2}}$
+$=\frac{1}{\phi^{2}}\sum_{i=1}^{n}w_{i}^{2}\frac{\tfrac{\phi}{w_{i}}\zeta_{i}^{p}}{\zeta_{i}^{2p-2}}x_{i,j_{1}}x_{i,j_{2}}$
+$=\frac{1}{\phi}\sum_{i=1}^{n}w_{i}\zeta_{i}^{2-p}x_{i,j_{1}}x_{i,j_{2}}$
+$=\frac{1}{\phi}\bm{X}^{\top}\bm{D}(W)\bm{X}$
+$=-A.$
+Hence, B=AB=-A, and the asymptotic covariance matrix reduces to
+$\Sigma=-A^{-1}.$